### v6

# Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions

Abstract
In this paper, we investigate the scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes. These processes exhibit a power-law tail of the form $n^{-(1+\alpha)}$ for $\alpha\in(\frac{1}{2},1)$, and the $\ell^{1}$ norm of the kernel vector converges to 1. We demonstrate that the discrete-time scaling limit retains a long-memory property and can be viewed as an integrated fractional Cox-Ingersoll-Ross process. Moreover, we present an efficient method for simulating the fractional Cox-Ingersoll-Ross process. The simulation and Goodness-of-Fit Test code are available at https://github.com/gagawjbytw/INAR-rough-Heston.
Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions
Yingli Wang  [^1] School of Mathematics, Shanghai University of Finance and Economics, Shanghai, People’s Republic of China; 2022310119@163.sufe.edu.cn., Chunhao Cai  [^2] School of Mathematics (Zhuhai), Sun Yat-Sen University, Zhuhai, People’s Republic of China; caichh9@mail.sysu.edu.cn, Ping He  [^3] School of Mathematics, Shanghai University of Finance and Economics, Shanghai, People’s Republic of China; pinghe@mail.shufe.edu.cn, Qinghua Wang  [^4] School of Mathematics, Shanghai University of Finance and Economics, Shanghai, People’s Republic of China; wangqinghua@mail.shufe.edu.cn
Keywords: heavy tailed unstable INAR($\infty$) processes, fractional CIR processes, rough volatility, option pricing.
MSC Classification: 60G22, 60F05
1 Introduction
Let $(\epsilon_{n})_{n\in\mathbb{Z}}$ be i.i.d. Poisson($\nu$) random variables. Let $(\eta_{k})_{k\geq 1}$ be a sequence of positive numbers in $\ell^{1}$. For $n\in\mathbb{Z}$, $k\geq 1$, and $l\geq 1$, let $\xi_{l}^{(n,k)}$ be mutually independent random variables with a Poisson($\eta_{k}$) distribution, which are also independent of the sequence $(\epsilon_{n})_{n\in\mathbb{Z}}$.
An Integer-Valued Autoregressive process of infinite order, abbreviated as INAR($\infty$), refers to a type of integer-valued time series model denoted by $(X_{n})_{n\in\mathbb{Z}}$. This model constitutes an infinite-order extension of the classic INAR($p$) processes. The process is formally defined by the following system of stochastic difference equations:
$\epsilon_{n}=X_{n}-\sum_{k=1}^{\infty}\sum_{l=1}^{X_{n-k}}\xi_{l}^{(n,k)},\quad n\in\mathbb{Z}.$
To establish a more compact notation, we introduce the reproduction operator, denoted by $\circ$”. For a reproduction parameter $\alpha\geq 0$, we define the offspring sequence $(\xi_{n}^{(\alpha)})_{n\geq 1}$ as a sequence of i.i.d. random variables, where each element $\xi_{n}^{(\alpha)}$ is an offspring variable following a Poisson distribution with mean $\alpha$.
The reproduction operator is then defined for any non-negative integer-valued random variable $Y$ as the sum of its offspring:
$\alpha\circ Y:=\sum_{n=1}^{Y}\xi_{n}^{(\alpha)}$
provided that $Y$ is independent of the offspring sequence $(\xi_{n}^{(\alpha)})$. Using this operator, the system of stochastic difference equations can be compactly rewritten as:
$\epsilon_{n}=X_{n}-\sum_{k=1}^{\infty}\eta_{k}\circ X_{n-k}.$
Finally, we define $\nu$ as the immigration parameter, $(\epsilon_{n})$ as the immigration sequence, and $\eta_{k}$ as the reproduction coefficient for each non-negative integer $k$.
A cumulative INAR($\infty$) process (also known as a discrete Hawkes process) , which starts from time $1$ and is denoted by $(N_{n})_{n\geq 1}$, is defined by
$N_{n}=\sum_{s=1}^{n}X_{s}.$ (1)
Hawkes processes, introduced by Hawkes (1971), are continuous-time self-exciting point processes widely used in various fields. Discrete-time analogs, such as cumulative INAR($\infty$) processes, provide similar modeling capabilities, particularly for count data observed at fixed time intervals. However, the scaling limits of heavy-tailed, unstable cumulative INAR($\infty$) processes differ considerably from those of continuous-time counterparts, which warrants a dedicated investigation. Under certain conditions, the Poisson autoregressive process can be viewed as an INAR($\infty$) process with Poisson offspring. For a comprehensive discussion on Poisson autoregressive models and their connections to INAR and Hawkes processes, refer to Fokianos (2024), Huang and Khabou (2023) and Kirchner (2016). From Proposition 3 in Kirchner (2016), we see that if an INAR($\infty$) process $(X_{n})_{n\geq 1}$ starts from time $1$, it can also be defined by:
$\lambda_{n}:=\nu+\sum_{s=1}^{n-1}\eta_{n-s}X_{s},$ (2)
where $\nu>0$ is the immigration rate, and $(\eta_{n})_{n\geq 1}\in\ell^{1}$ represents the offspring distribution with $\eta_{n}\geq 0$ for all $n\in\mathbb{N}$. Given the history $\mathcal{F}_{n-1}$, the count $X_{n}$ follows a Poisson distribution with parameter $\lambda_{n}$, i.e.,
$X_{n}\mid\mathcal{F}_{n-1}\sim\text{Poisson}(\lambda_{n}).$
Scaling limits for self-exciting processes, a topic of significant theoretical and applied interest, have been extensively studied. For example, consider a Hawkes process $(N_{t})_{t\geq 0}$ with intensity function $h$. A common goal is to characterize the asymptotic behavior of the rescaled process
$(\beta_{T}N_{tT})_{t\in[0,1]},$
as $T\rightarrow\infty$, where $\beta_{T}$ is an appropriate normalizing factor. In Bacry et al. (2013), it is shown that under the condition
$\left\lVert h\right\rVert_{L^{1}}\equiv\int_{0}^{\infty}h(s)ds<1,$
the asymptotic behavior of a Hawkes process is quite similar to that of a Poisson process. Indeed, as $T$ tends to infinity,
$\sup_{t\in[0,1]}\left|\frac{N_{tT}}{T}-\mathbb{E}\left[\frac{N_{tT}}{T}\right]\right|\rightarrow 0\ \text{in probability},$
and
$\left(\sqrt{T}\left(\frac{N_{tT}}{T}-\mathbb{E}\left[\frac{N_{tT}}{T}\right]\right)\right)_{t\in[0,1]}\rightarrow\sigma(W_{t})_{t\in[0,1]}\ \text{in law}$
under the skorohod topology, with $\sigma$ an explicit constant and $(W_{t})$ a Brownian motion. This result has been extended in Zhu (2013) to the case of nonlinear Hawkes processes. Regarding unstable Hawkes processes, the light-tailed case and the heavy-tailed case are respectively studied in Jaisson and Rosenbaum (2015) and Jaisson and Rosenbaum (2016). The light-tailed case corresponds to a CIR limit obtained under the crucial assumption
$\int_{0}^{\infty}sh(s)ds<+\infty$
and the heavy-tailed case corresponds to a fractional CIR limit obtained under
$\int_{0}^{\infty}sh(s)ds=+\infty\ \text{and}\ h(x)\sim\frac{K}{x^{1+\alpha}}\ \text{as}\ x\rightarrow\infty.$
In the fields of mathematical finance and econometrics, for example, the limit behavior of self-exciting dynamics is of particular interest. Studies such as Horst and Xu (2019, 2022) examine the microstructure of financial markets and the impact of external shocks on market dynamics. They mathematically demonstrate that multiple small external shocks can trigger endogenous jump cascades in asset returns and stock price volatility.
When it comes to the discrete case, the asymptotic behavior of unstable Integer-Valued Autoregressive processes of order $p$, or INAR($p$), was investigated by Barczy et al. (2011). An INAR($p$) process $(X_{n})$ is defined by the stochastic difference equation
$X_{n}=\sum_{i=1}^{p}\alpha_{i}\circ X_{n-i}+\epsilon_{n},$
where the $\alpha_{i}\in[0,1)$ are autoregressive parameters, $\circ$ denotes the binomial thinning operator, and $(\epsilon_{n})$ is a sequence of i.i.d. non-negative integer-valued random variables representing an immigration term. The authors proved that under the critical condition where the sum of these autoregressive parameters equals one:
$\sum_{i=1}^{p}\alpha_{i}=1,$
a sequence of appropriately scaled random step functions formed from the process converges weakly towards a squared Bessel process. This result can be extended to INAR($\infty$) processes, which corresponds to what we term the “light-tailed” case. Consequently, the heavy-tailed case has not been addressed in the literature, a gap this paper aims to fill.
In this paper, we aim to provide a discrete-time counterpart to the continuous-time results found in Jaisson and Rosenbaum (2016), focusing on heavy-tailed unstable INAR($\infty$) processes. Specifically, we investigate an INAR($\infty$) process with
$\eta_{n}\sim\frac{K}{n^{1+\alpha}}\quad\text{as}\quad n\to\infty,$
where the constant $\alpha\in\left(\frac{1}{2},1\right)$. Inspired by the continuous-time case in Jaisson and Rosenbaum (2016), we consider a sequence of INAR processes $(X_{n}^{T})_{n\in\mathbb{N}}$ and analyze the asymptotic behavior as $T\to\infty$. The sequence of parameters $(\lambda_{n}^{T})$ is defined by
$\lambda_{n}^{T}=\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}X_{s}^{T},\quad n\in\mathbb{N},$ (3)
where $\mu^{T}$ is a sequence of positive real numbers. As in Jaisson and Rosenbaum (2016), we establish specific assumptions on $\eta^{T}$ to facilitate our analysis. We establish the tightness of the renormalized cumulative INAR($\infty$) process $Y^{T}$ and associated martingale $Z^{T}$ defined in (5) and (7) which will be given below under specific assumptions, and show that any limit point $(Z,Y)$ of $(Z^{T},Y^{T})$ converges to a rough fractional diffusion process characterized by an integrated fractional Cox-Ingersoll-Ross (CIR) process.
While our work is inspired by and presents a discrete-time counterpart to the celebrated results of Jaisson and Rosenbaum (2016), it is motivated by distinct theoretical and practical considerations that underscore its unique contribution. The transition from a continuous-time Hawkes process to a discrete-time INAR($\infty$) process is far from trivial and offers several key advantages:
  1. 1.
    Direct Modeling of Intrinsically Discrete Data: Many real-world phenomena, particularly in finance (e.g., tick-by-tick trade counts, order book events) and social sciences (e.g., daily event counts), are recorded as discrete counts over fixed time intervals. The INAR($\infty$) framework provides a more natural and direct modeling tool for such data, avoiding the potentially lossy abstraction required to fit them into a continuous-time point process framework.
  2. 2.
    Novel Proof Techniques and Theoretical Insights: The proofs in the discrete setting are not mere translations of their continuous-time counterparts. They require different mathematical machinery, such as discrete renewal theory and specific martingale central limit theorems for discrete processes. This distinct theoretical path reveals unique structural properties of the discrete models, such as the specific long-memory characteristics we highlight.
  3. 3.
    A Computationally Superior Simulation Framework: Perhaps the most significant practical contribution is that INAR($\infty$) processes are substantially easier and more efficient to simulate than their continuous-time, singular-kernel Hawkes counterparts. Simulating our model only requires iterative sampling from a Poisson distribution. Our scaling limit theorem thus serves as a crucial bridge: it rigorously validates that this simple, efficient discrete simulation converges to the complex continuous process of interest. This directly addresses the question of “why not just use a Hawkes process?”: The INAR($\infty$) process provides a computationally tractable pathway for its analysis and application.
  4. 4.
    New Applications in Financial Modeling: This framework opens the door to new applications not explored in the original Hawkes process literature. For instance, as demonstrated in the recent work of Wang and Cui (2025), a bivariate version of this INAR($\infty$) framework can be shown to converge to the rough Heston model. This provides not only a novel microstructural foundation for rough volatility but also a highly efficient simulation engine for pricing both European and path-dependent options, a task that remains challenging for other methods.
Our primary result, encapsulated in Theorem 1, demonstrates that under heavy-tailed and unstable conditions, the cumulative INAR($\infty$) process converges to a rough fractional diffusion process. This finding is achieved by leveraging discrete renewal theory and martingale central limit theorems.
These results contribute to the theoretical understanding of discrete self-exciting processes by revealing inherent long-memory characteristics under heavy-tailed conditions. Such insights have significant implications for fields such as finance, epidemiology, and network traffic analysis, where capturing long-range dependencies and extreme events is crucial for accurate modeling and forecasting.
2 Assumptions and Main Results
First, we define an unstable cumulative INAR process with $\|\eta\|_{1}=1$. We propose two assumptions similar to those in Jaisson and Rosenbaum (2016).
Assumption 1.
We assume the following conditions hold:
  1. 1.
    The baseline offspring distribution $(\eta_{n})_{n\geq 1}$ has a heavy tail, characterized by the asymptotic behavior:
    $\eta_{n}\sim\frac{K}{n^{1+\alpha}}\quad\text{as}\quad n\to\infty,$
    for some $\alpha\in(\frac{1}{2},1)$ and some positive constant $K$.
  2. 2.
    The sequence of scaling factors $(a_{T})_{T\geq 0}$ consists of positive numbers such that for all $T$, $a_{T}<1$, and it converges to 1 as $T\to\infty$.
  3. 3.
    For each $T$, the reproduction coefficients are given by $\eta_{n}^{T}=a_{T}\eta_{n}$ for all $n\geq 1$.
The assumption on the asymptotic behavior of $\eta_{n}$ in Assumption 1 is a standard condition for regularly varying sequences. A direct consequence of this assumption, based on Karamata’s Tauberian Theorem for power series, is that the tail sum of the sequence also has a well-defined asymptotic form. Specifically, it implies that:
$\lim_{n\to\infty}\alpha n^{\alpha}\sum_{k=n+1}^{\infty}\eta_{k}=K.$
This equivalent formulation of the tail behavior is what we primarily use in our subsequent proofs.
Assumption 2.
There exist two positive constants $\lambda,\nu^{*}$ such that
$\lim_{T\rightarrow\infty}T^{\alpha}(1-a_{T})=\lambda\delta,$
and
$\lim_{T\rightarrow\infty}T^{1-\alpha}\mu^{T}=\nu^{*}\delta^{-1},$
where $\delta=K\frac{\Gamma(1-\alpha)}{\alpha}$. The sequence $(\mu^{T})$ is shown in (3).
We first introduce some notation. Define
$M_{n}^{T}:=\sum_{s=1}^{n}\left(X_{s}^{T}-\lambda_{s}^{T}\right)=N_{n}^{T}-\sum_{s=1}^{n}\lambda_{s}^{T}.$
We also set $(A_{n}^{T})_{n\geq 1}\in\ell^{1}$ as the sequence defined on $\mathbb{N}$ by
$A_{n}^{T}:=\sum_{k=1}^{\infty}(\eta^{T})_{n}^{*k},$ (4)
where * denotes the discrete convolution, which, for two non-negative sequences $(q_{n})_{n\geq 1}$, $(m_{n})_{n\geq 1}\in\ell^{1}$,
$(q*m)(n)=\sum_{s=1}^{n-1}q_{s}m_{n-s},$
and $\eta^{*k+1}$ denotes the discrete convolution of $\eta^{*k}$ with $\eta$, i.e. $\eta^{*k+1}=\eta*\eta^{*k}$. $(A_{n}^{T})_{n\geq 1}$ is well-defined since $\left\lVert\eta^{T}\right\rVert_{1}<1$.
Let us introduce our main result. We consider for $t\in[0,1]$ the renormalized cumulative INAR($\infty$) process
$Y_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=1}^{\lfloor Tt\rfloor}X_{s}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}N_{\lfloor Tt\rfloor}^{T},$ (5)
and its associated cumulative intensity
$\Lambda_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=1}^{\lfloor Tt\rfloor}\lambda_{s}^{T}.$ (6)
The martingale $Z^{T}$ is defined as
$Z_{t}^{T}:=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}(Y_{t}^{T}-\Lambda_{t}^{T}).$ (7)
Recall the definition of the function $f^{\alpha,\lambda}$ from Section 3.1 in Jaisson and Rosenbaum (2016), which is
$f^{\alpha,\lambda}(x)=\lambda x^{\alpha-1}E_{\alpha,\alpha}(-\lambda x^{\alpha})$ (8)
and
$F^{\alpha,\lambda}(t)=\int_{0}^{t}f^{\alpha,\lambda}(x)dx,$ (9)
where $E_{\alpha,\beta}$ is the Mittag-Leffler function, defined as
$E_{\alpha,\beta}(z)=\sum_{n=0}^{\infty}\frac{z^{n}}{\Gamma(\alpha n+\beta)}.$
For a comprehensive review of its properties and applications, see, e.g., Haubold et al. (2011). We are now ready to state our results about the convergence in distribution of $(Z^{T},Y^{T})$ under the skorohod topology. We first introduce the following lemma, a discrete version of Lemma 4.3 in Jaisson and Rosenbaum (2016), which will be useful in the proof of equation (12) and Proposition 1.
Lemma 1.
Define a sequence of functions $(\rho^{T})_{T\geq 0}$ on $0,\infty)$ by
$\rho^{T}(x):=T\frac{A_{\lfloor xT\rfloor}^{T}}{\left\lVert A^{T}\right\rVert_{1}}.$ (10)
Each $\rho^{T}$ is a step function and a valid probability density function, since it is non-negative and integrates to one. As $T\to\infty$, the sequence of probability measures corresponding to these densities converges weakly to a probability measure on $0,\infty)$ with the density $\lambda x^{\alpha-1}E_{\alpha,\alpha}(-\lambda x^{\alpha})$.
Furthermore, let $F^{T}(t)$ be the associated cumulative distribution function, defined by $F^{T}(t)=\int_{0}^{t}\rho^{T}(x)dx$. Then, over the compact interval $[0,1]$, $F^{T}(t)$ converges uniformly to the limiting distribution function $F^{\alpha,\lambda}(t)$ defined in (9).
Proof.
We provide the proof in Section 4.1. ∎
To prove this lemma, we need a discrete-type extended Karamata-Tauberian Theorem, which is similar to the Theorem 1.7.6 in Bingham et al. (1989). We first give the discrete versions of Proposition 1.5.8, Theorem 1.7.6 in Bingham et al. (1989).
Lemma 2.
Let $l(t)$ be a slowly varying function that is locally bounded on $X_{0},\infty)$ for some $X_{0}>0$, and let $\alpha>-1$. For any fixed integer $X\geq X_{0}$, the following holds for integers $x\to\infty$:
$\sum_{t=X}^{x}t^{\alpha}l(t)\sim\frac{1}{\alpha+1}x^{\alpha+1}l(x).$
Proof.
We provide the proof in Section 4.2. ∎
Lemma 3 (Discrete-type extended Karamata-Tauberian Theorem).
Let $(\beta_{n})_{n\geq 0}$ be a non-negative sequence, $c\geq 0$, $\rho>-1$, and let $l$ be a slowly varying function. For $s>0$, define the transform
$\hat{\beta}(s):=\sum_{n=0}^{\infty}(e^{-sn}-e^{-s(n+1)})\beta_{n}=(1-e^{-s})\sum_{n=0}^{\infty}e^{-sn}\beta_{n}.$
Assume this series converges for all $s>0$. If
$\beta_{n}\sim\frac{cn^{\rho}l(n)}{\Gamma(1+\rho)}\quad\text{as }n\to\infty,$
then
$\hat{\beta}(s)\sim cs^{-\rho}l(1/s)\quad\text{as }s\downarrow 0.$
Proof.
We provide the proof in Section 4.3. ∎
The tightness of $(Z^{T},Y^{T})$ also requires proof. Therefore, we start with the following lemma.
Lemma 4.
The sequences $(Y^{T})$ and $(\Lambda^{T})$ are $C$-tight.
Proof.
We provide the proof in Section 4.4. ∎
Since $(Y^{T})$ and $(\Lambda^{T})$ are pure jump processes, the classical Kolmogorov-Chentsov tightness criterion is not applicable. To establish the tightness of $(Z^{T},Y^{T})$, we instead apply a recent criterion from Horst et al. (2023), which we restate as Lemma 5 below for the reader’s convenience. This result gives a new criterion to verify the $C$-tightness for a sequence of càdlàg processes and will be used in proving Lemma 4. Additionally, proving the convergence of the sharp bracket of $Z^{T}$ in the discrete-time case presents more challenges than in the continuous-time case discussed by Jaisson and Rosenbaum (2016). For instance, in a Hawkes process $(N_{t})_{t\geq 0}$ with intensity function is
$\lambda_{t}=\nu+\int_{0}^{t}h(t-s)\,dN_{s},$
the angle bracket of $(N_{t})_{t\geq 0}$ is given by $\int_{0}^{\cdot}\lambda_{s}\,ds$, and the sharp bracket is simply $N_{\cdot}$, which is straightforward to derive. However, achieving a similar result in the discrete-time setting is not as immediate and requires more nuanced handling. Indeed, the value $\lambda_{n}$ in (2) is analogous the intensity process $\lambda_{t}$ in continuous-time Hawkes processes. In the proof section, we work directly with $(Y^{T})$, and in Section 2.1 provides an intuitive interpretation of the rescaled $\lambda^{T}$. To clarify the proof clear, we introduce following lemma here.
Lemma 5 (Criterion for $C$-tightness, Lemma 3.5 in Horst et al. (2023)).
Let $(X^{(n)})_{n\geq 1}$ be a sequence of càdlàg stochastic processes, where each process $X^{(n)}=(X_{t}^{(n)})_{t\geq 0}$ is defined on a suitable probability space. Let $\Delta_{h}X_{t}^{(n)}:=X_{t+h}^{(n)}-X_{t}^{(n)}$ denote the increment of the process $X^{(n)}$ over a time interval of length $h$. If for some $T>0$ the following three conditions hold:
  1. 1.
    $\sup_{n\geq 1}\mathbb{E}\left[\left|X^{(n)}_{0}\right|^{\beta}\right]<\infty$
    for some $\beta>0$;
  2. 2.
    $\sup_{k=0,1,\cdots,\lfloor Tn^{\theta}\rfloor}\sup_{h\in[0,1/n^{\theta}]}|\Delta_{h}X_{k/n^{\theta}}^{(n)}|\rightarrow 0\ \text{in probability as}\ n\rightarrow\infty,$
    for some $\theta>2$;
  3. 3.
    there exist some constants $C>0$, $p\geq 1$, $m\in\{1,2,\cdots\}$, $\theta>2$ and pairs $\{(a_{i},b_{i})\}_{i=1,\cdots,m}$ satisfying
    $a_{i}\geq 0,\ b_{i}>0,\ \min_{1\leq i\leq m}\{b_{i}+a_{i}/\theta\}>1,$
    such that for all $n\geq 1$ and $h\in(0,1)$,
    $\sup_{t\in[0,T]}\mathbb{E}\left[\left|\Delta_{h}X_{t}^{(n)}\right|^{p}\right]\leq C\cdot\sum_{i=1}^{m}\frac{h^{b_{i}}}{n^{a_{i}}}.$
Then the sequence $(X^{(n)})_{n\geq 1}$ is $C$-tight on $[0,T]$ (i.e., tight in the space of continuous functions $C([0,T])$ equipped with the uniform topology).
Additionally, we need the following lemma to prove tightness and convergence in the Skorohod topology.
Lemma 6.
The sequence of martingales $Y^{T}-\Lambda^{T}$ tends to zero in probability, uniformly on $[0,1]$.
Proof.
We provide the proof in Section 4.5. ∎
Now we can give the main result.
Proposition 1 (Tightness of the Sequence).
Under Assumption 1 and 2, the sequence $(Z^{T},Y^{T})$ is tight. Furthermore, if $(Z,Y)$ is a limit point of $(Z^{T},Y^{T})$ as $T\rightarrow\infty$, then $Z$ is a continuous martingale and $[Z,Z]=Y$.
Proof.
We provide the proof in Section 5.1. ∎
Next, we characterize the limit process $Y$ (see Theorem 1). To this end, we first introduce a lemma that solves discrete renewal equations, which will be essential for proving Theorem 1.
Lemma 7 (Solution of Discrete Renewal Equations).
Given a non-negative sequence $(\eta_{n})_{n\geq 1}\in\ell^{1}$ and two non-negative sequences $(x_{n})_{n\geq 1}$, $(y_{n})_{n\geq 1}$, the following equation
$x_{n}=y_{n}+\sum_{s=1}^{n-1}\eta_{s}x_{n-s}$ (11)
has the unique solution
$x_{n}=\left(y+y*A\right)(n)=y_{n}+\sum_{i=1}^{n-1}A_{i}y_{n-i},$
where $(A_{n})_{n\geq 1}$ is defined in (4).
Proof.
We provide the proof in Section 4.6. ∎
Theorem 1 (Characterization of the Limit Process).
For any limit point $(Z,Y)$ defined in Proposition 1, there exists a limiting probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which the convergence holds almost surely. On this probability space, there exists a Brownian motion $B$ (up to extension) such that for all $t\in[0,1]$, $Z_{t}=B_{Y_{t}}$. Furthermore, for any $\epsilon>0$, $Y$ is continuous with Hölder regularity $(1\wedge 2\alpha)-\epsilon$ on $[0,1]$ and satisfies
$Y_{t}=\int_{0}^{t}sf^{\alpha,\lambda}(t-s)\,ds+\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)B_{Y_{s}}\,ds.$ (12)
Proof.
The proof of the Hölder property for $Y$ in Theorem 1 follows exactly the same argument as the Proof of Theorem 3.1 in Jaisson and Rosenbaum (2016). The derivation of the equation (12) is provided in Section 5.2. ∎
After (12) in Theorem 1 is established, we can characterize its derivative process as follows. This result is from Theorem 3.2 in Jaisson and Rosenbaum (2016).
Corollary 1 (Characterization of the Derivative Process).
Let $(Y_{t})$ be a process satisfying (12) for $t\in[0,1]$ and assume $\alpha>1/2$. Then $Y$ is differentiable on $[0,1]$ and its derivative $\dot{Y}$ satisfies
$\dot{Y}_{t}=F^{\alpha,\lambda}(t)+\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)\sqrt{\dot{Y}_{s}}dB_{s},$
where $B$ is the same Brownian motion as in (12). Furthermore, for any $\epsilon>0$, $\dot{Y}$ has Hölder regularity $\alpha-1/2-\epsilon$.
Proof.
The proof is exactly the same as the Proof of Theorem 3.2 in Jaisson and Rosenbaum (2016). ∎
2.1 Heuristic Motivation for the Scaling Limit
In the “light-tailed” case, when the condition $\sum_{n=1}^{\infty}n\eta_{n}<\infty$ is met as in Barczy et al. (2011), the scaling limit of the INAR process is known to be a squared Bessel process. Our work, however, focuses on the heavy-tailed case governed by Assumption 1, which induces a different scaling behavior.
To gain an intuitive understanding of this limit, we follow an approach inspired by the analysis of continuous-time Hawkes processes by Jaisson and Rosenbaum (2016). The central idea is to study the asymptotic behavior of a properly renormalized version of the conditional intensity process, $(\lambda_{n}^{T})$. Let us define the renormalized intensity process $(C_{t}^{T})_{t\in[0,1]}$ as:
$C_{t}^{T}:=\frac{1-a_{T}}{\mu^{T}}\lambda_{\lfloor tT\rfloor}^{T}.$ (13)
The subsequent analysis in this section will heuristically show how this discrete process can be approximated by a continuous-time fractional Cox-Ingersoll-Ross process as $T\to\infty$. This provides an intuitive justification for the main results presented in our theorems.
To begin this heuristic analysis, we first introduce an alternative expression for $\lambda_{n}^{T}$ that will be useful in our derivations.
Lemma 8.
For every $n\in\mathbb{N}$,
$\lambda_{n}^{T}=\mu^{T}+\mu^{T}\sum_{s=1}^{n-1}A_{n-s}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right).$ (14)
Proof.
We provide the proof in Section 4.7. ∎
We substitute the expression for $\lambda_{\lfloor tT\rfloor}^{T}$ from Lemma 8 to expand the definition of $C_{t}^{T}$:
$C_{t}^{T}=$ $\frac{(1-a_{T})}{\mu^{T}}\lambda_{\lfloor tT\rfloor}^{T}=\frac{(1-a_{T})}{\mu^{T}}\left(\mu^{T}+\mu^{T}\sum_{s=1}^{\lfloor tT\rfloor-1}A_{s}^{T}+\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor tT\rfloor-s}^{T}(M_{s}^{T}-M_{s-1}^{T})\right)$ (15)
$=$ $(1-a_{T})+(1-a_{T})\sum_{s=1}^{\lfloor tT\rfloor-1}A_{s}^{T}+\frac{(1-a_{T})}{\mu^{T}}\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor tT\rfloor-s}^{T}(M_{s}^{T}-M_{s-1}^{T}).$
The asymptotic form of the renewal sequence $A_{k}^{T}$ was established in the proof of Lemma 1. For large $T$ and $k$, we have the approximation:
$A_{k}^{T}\approx\frac{\left\lVert A^{T}\right\rVert_{1}v_{T}}{T}\left(\frac{k}{T}\right)^{\alpha-1}E_{\alpha,\alpha}\left(-v_{T}\left(\frac{k}{T}\right)^{\alpha}\right).$
Using this approximation, the sum can be viewed as a Riemann sum, which for large $T$ converges to the corresponding integral:
$\sum_{s=1}^{\lfloor tT\rfloor-1}A_{s}^{T}$ $=\sum_{s=1}^{\lfloor tT\rfloor-1}\frac{1}{T}\cdot\left(TA_{s}^{T}\right)$
$\sim\int_{0}^{t}TA_{\lfloor Tx\rfloor}^{T}dx$
$\sim\int_{0}^{t}\left\lVert A^{T}\right\rVert_{1}v_{T}x^{\alpha-1}E_{\alpha,\alpha}(-v_{T}x^{\alpha})dx,$
where $v_{T}$ is defined in (23). Define
$B_{t}^{T}:=\frac{1}{\sqrt{T}}\sum_{s=1}^{\lfloor tT\rfloor-1}\frac{M_{s}^{T}-M_{s-1}^{T}}{\sqrt{\lambda_{s}^{T}}}.$
$(M_{s}^{T}-M_{s-1}^{T})_{s\geq 1}$ is a martingale difference sequence, and it is obvious that
$\sum_{s=1}^{\lfloor tT\rfloor-1}\mathbb{E}\left[\frac{1}{T}\frac{(M_{s}^{T}-M_{s-1}^{T})^{2}}{\lambda_{s}^{T}}\right]=$ $\frac{\lfloor tT\rfloor-1}{T}\rightarrow t,\ \text{as}\ T\rightarrow\infty.$
We also need to check the conditional Lindeberg condition for the martingale central limit theorem. Let the martingale differences be $\Delta B_{s}^{T}=\frac{1}{\sqrt{T}}\frac{X_{s}^{T}-\lambda_{s}^{T}}{\sqrt{\lambda_{s}^{T}}}$. The condition requires that for any $\epsilon>0$, the sum of conditional expectations,
$L_{T}:=\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}\left[(\Delta B_{s}^{T})^{2}\cdot 1_{\{|\Delta B_{s}^{T}|>\epsilon\}}\bigg|\mathcal{F}_{s-1}\right],$
converges to 0 in probability as $T\to\infty$.
To prove this, we first find a bound for each term in the sum using a standard moment inequality. Given the history $\mathcal{F}_{s-1}$, $\lambda_{s}^{T}$ is a known quantity. For the random variable $\Delta B_{s}^{T}$ and constant $\epsilon>0$, we have:
$\mathbb{E}\left[(\Delta B_{s}^{T})^{2}\cdot 1_{\{|\Delta B_{s}^{T}|>\epsilon\}}\bigg|\mathcal{F}_{s-1}\right]$ $\leq\frac{\mathbb{E}\left[(\Delta B_{s}^{T})^{4}\bigg|\mathcal{F}_{s-1}\right]}{\epsilon^{2}}$
$=\frac{1}{\epsilon^{2}T^{2}}\mathbb{E}\left[\left.\frac{(X_{s}^{T}-\lambda_{s}^{T})^{4}}{(\lambda_{s}^{T})^{2}}\right|\mathcal{F}_{s-1}\right]$
[…]
$=\frac{1}{\epsilon^{2}T^{2}}\left(\frac{1}{\lambda_{s}^{T}}+3\right).$
Summing this bound from $s=1$ to $\lfloor tT\rfloor$ provides a random upper bound for $L_{T}$:
$0\leq L_{T}\leq\frac{1}{\epsilon^{2}T^{2}}\sum_{s=1}^{\lfloor tT\rfloor}\left(\frac{1}{\lambda_{s}^{T}}+3\right)=:R_{T}.$
To show that $L_{T}\to 0$ in probability, it is sufficient to show that $R_{T}\to 0$ in probability. We prove this by showing that its expectation converges to zero, i.e., $\mathbb{E}[R_{T}]\to 0$.
$\mathbb{E}[R_{T}]$ $=\mathbb{E}\left[\frac{1}{\epsilon^{2}T^{2}}\sum_{s=1}^{\lfloor tT\rfloor}\left(\frac{1}{\lambda_{s}^{T}}+3\right)\right]=\frac{1}{\epsilon^{2}T^{2}}\sum_{s=1}^{\lfloor tT\rfloor}\left(3+\mathbb{E}\left[\frac{1}{\lambda_{s}^{T}}\right]\right).$
Since the process intensity is always bounded below by the baseline rate, $\lambda_{s}^{T}\geq\mu^{T}>0$, we have a bound for the expectation:
$\mathbb{E}\left[\frac{1}{\lambda_{s}^{T}}\right]\leq\frac{1}{\mu^{T}}.$
Substituting this into the expression for $\mathbb{E}[R_{T}]$:
$\mathbb{E}[R_{T}]$ $\leq\frac{1}{\epsilon^{2}T^{2}}\sum_{s=1}^{\lfloor tT\rfloor}\left(3+\frac{1}{\mu^{T}}\right)=\frac{\lfloor tT\rfloor}{\epsilon^{2}T^{2}}\left(3+\frac{1}{\mu^{T}}\right).$
From Assumption 2, we know $\mu^{T}\sim CT^{\alpha-1}$ for some constant $C>0$. Since $\alpha>1/2$, we have $\mathbb{E}[R_{T}]\to 0$ as $T\to\infty$. As $R_{T}$ is a non-negative random variable, convergence of its expectation to zero implies that $R_{T}\to 0$ in probability. As a result, $L_{T}$ must also converge to 0 in probability, which rigorously verifies the conditional Lindeberg condition.
By the martingale central limit theorem (see Section 3 in Hall and Heyde (1980)), $B_{t}^{T}$ converges to a standard Brownian motion. For sufficiently large $T$, from Lemma 1,
$\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor tT\rfloor-s}^{T}(M_{s}^{T}-M_{s-1}^{T})$ (16)
$=$ $\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor tT-s\rfloor}^{T}(M_{s}^{T}-M_{s-1}^{T})$
$=$ $\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor T\left(t-\frac{s}{T}\right)\rfloor}^{T}(M_{s}^{T}-M_{s-1}^{T})$
$\approx$ $\sum_{s=1}^{\lfloor tT\rfloor-1}\frac{\left\lVert A^{T}\right\rVert_{1}}{T}v_{T}\cdot\left(t-\frac{s}{T}\right)^{\alpha-1}E_{\alpha,\alpha}\left(-v_{T}\left(t-\frac{s}{T}\right)^{\alpha}\right)\sqrt{T\lambda_{s}^{T}}\left(B_{(s+1)/T}^{T}-B_{s/T}^{T}\right).$
So here we face the same problem as in Jaisson and Rosenbaum (2016): to ensure that the limits of $v_{T}$ and another related term $\gamma_{T}$ are positive constants. This motivates our Assumption 2, which is consistent with the framework in Jaisson and Rosenbaum (2016).
Under Assumption 2, let
$\kappa_{s}^{T}=\sum_{u=1}^{s-1}A_{s-u}^{T}(M_{u}^{T}-M_{u-1}^{T}),$
and
$\gamma_{T}=\frac{1}{\sqrt{\mu^{T}T(1-a_{T})}},$
then (16) can be written as
$\frac{1}{\left\lVert A^{T}\right\rVert_{1}\mu^{T}}\kappa_{\lfloor sT\rfloor}^{T}\approx$ $\gamma_{T}v_{T}\int_{0}^{s}\left(s-u\right)^{\alpha-1}E_{\alpha,\alpha}\left(-v_{T}\left(s-u\right)^{\alpha}\right)\sqrt{C_{u}^{T}}dB_{u}^{T}$ (17)
$\stackrel{{\scriptstyle\text{w}}}{{\rightarrow}}$ $\gamma_{\infty}v_{\infty}\int_{0}^{s}(s-u)^{\alpha-1}E_{\alpha,\alpha}(-v_{\infty}(s-u)^{\alpha})\sqrt{C_{u}^{\infty}}dB_{u},$
as $T\rightarrow\infty$, where
$\gamma_{\infty}=\frac{1}{\sqrt{\lambda\nu^{*}}},v_{\infty}=\lambda.$
2.2 A Microstructural Foundation for the Initial Value
In practical applications, particularly for financial models like the rough Heston model, specifying the initial value of the volatility process, $\dot{Y}_{0}=v_{0}$, is essential. This is fundamentally different from setting the initial value of the integrated process, $Y_{0}$. Achieving this requires embedding the initial dynamics directly into the discrete-time microstructural model.
Following the methodology of El Euch and Rosenbaum (2019) and its discrete-time adaptation in Wang and Cui (2025), we introduce a carefully constructed time-dependent baseline intensity, $\hat{\mu}_{T}(n)$. This baseline is no longer a constant but is designed to create a burst of initial activity that, upon scaling, converges to the desired initial value $v_{0}$.
Let us redefine the intensity process from (3) in Section 2. The immigration component is now time-varying:
$\lambda_{n}^{T}=\hat{\mu}_{T}(n)+\sum_{s=1}^{n-1}\eta_{n-s}^{T}X_{s}^{T},\quad n\in\mathbb{N},$ (18)
where the time-dependent baseline intensity is given by:
Definition 1 (Time-Dependent Baseline for Initial Value).
For a given constant $\xi\geq 0$, the baseline intensity is defined as:
$\hat{\mu}_{T}(n):=\mu^{T}+\xi\mu^{T}\left(\frac{1}{1-a_{T}}\left(1-\sum_{s=1}^{n-1}\eta_{s}\right)-\sum_{s=1}^{n-1}\eta_{s}\right),$
where $(\eta_{n})_{n\geq 1}$ is the base kernel sequence (before scaling by $a_{T}$) satisfying $\sum_{n=1}^{\infty}\eta_{n}=1$. The parameter $\xi$ is a dimensionless constant that will directly control the initial value of the limit process.
The intuition for the above definition is as follows. At the very beginning, for $n=1$, the sums are empty, leading to a large initial baseline intensity:
$\hat{\mu}_{T}(1)=\mu^{T}\left(1+\frac{\xi}{1-a_{T}}\right).$
Since $(1-a_{T})^{-1}$ is of order $\mathcal{O}(T^{\alpha})$, this creates a significant initial burst of events. As $n$ increases, the term in the parenthesis decays, and $\hat{\mu}_{T}(n)$ quickly converges to the constant baseline $\mu^{T}$. This initial, engineered transient behavior is precisely what generates the non-zero starting value for the rate process $\dot{Y}_{t}$ in the continuous-time limit.
Now Theorem 1 can be extended to Corollary 2 which is given below. The Corollary shows that under this modified baseline intensity, the limit process $\dot{Y}_{t}$ starts from a value directly determined by $\xi$. We first introduce a Lemma which is useful in proving Corollary 2.
Lemma 9 (Renewal Expansion with Time-Dependent Baseline).
Under the time-dependent baseline intensity $\hat{\mu}_{T}(n)$ from Definition 1, the cumulative intensity and the centered cumulative process have the following representations for $n\geq 1$:
$\mathbb{E}[N_{n}^{T}]=\sum_{i=1}^{n}\left(\hat{\mu}_{T}(i)+\sum_{s=1}^{i-1}A_{i-s}^{T}\hat{\mu}_{T}(s)\right),$ (19)
and
$N_{n}^{T}-\mathbb{E}[N_{n}^{T}]=M_{n}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}M_{s}^{T}.$ (20)
Proof.
We provide the proof in Section 4.8. ∎
Corollary 2 (Characterization of the Derivative Process with Initial Process Value).
Under the assumptions of Theorem 1 and with the baseline intensity from Definition 1, the limit process $(\dot{Y}_{t})_{t\in[0,1]}$ is the unique solution to the stochastic Volterra equation
$\dot{Y}_{t}=F^{\alpha,\lambda}(t)+\xi\left(1-F^{\alpha,\lambda}(t)\right)+\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)\sqrt{\dot{Y}_{s}}dB_{s},$ (21)
which implies the initial condition $\dot{Y}_{0}=\xi$.
Proof.
We provide the proof in Section 5.3. ∎
This framework provides a theoretically sound and practical method for simulating the fractional CIR process from any desired starting level $v_{0}$ by directly encoding this value into the microscopic dynamics.
3 Numerical Validation and Application
The theoretical results of this paper show that the scaling limit of the heavy-tailed INAR($\infty$) process is an integrated fractional Cox-Ingersoll-Ross (CIR) process. This class of processes, governed by stochastic Volterra equations with singular kernels, has become central to modern financial mathematics for modeling so-called “rough volatility”, which captures the observed long-memory and low smoothness properties of market volatility time series.
However, the direct numerical simulation of such continuous-time fractional processes presents significant challenges. Standard discretization methods, such as the Euler-Maruyama scheme, often suffer from slow convergence rates and require extremely small time steps to achieve reasonable accuracy, making them computationally expensive.
3.1 Simulation via INAR($\infty$) Approximation
Our main theoretical result, encapsulated in Theorem 1 and Corollary 2, provides a novel and efficient pathway for simulating the fractional CIR process. The core idea is that we can generate sample paths of the limiting continuous-time process by simulating its discrete-time antecedent—the heavy-tailed cumulative INAR($\infty$) process—for a large but finite $T_{\text{steps}}$. This approach leverages the computational simplicity of simulating discrete count processes, which only requires iterative sampling from a Poisson distribution.
A key feature of our method is the ability to specify a non-zero initial value, $v_{0}$, for the fractional CIR process ($\dot{Y}_{t}$). This is not achieved by an ad-hoc adjustment post-simulation, but is instead built directly into the microstructural dynamics of the INAR($\infty$) process. As outlined in Definition 1, we employ a time-dependent baseline intensity, $\hat{\mu}_{T}(n)$, where the parameter $\xi$ is set to the desired initial value $v_{0}$. This creates an initial burst of activity that correctly sets the starting point of the process in the continuous-time limit.
The complete simulation procedure is detailed in Algorithm 1.
Algorithm 1 Simulation of a fractional CIR process via the cumulative INAR($\infty$) approximation
1:Input: Target initial value $v_{0}$, parameters $\alpha$, $\lambda_{\text{param}}$, $\nu^{*}$, and number of time steps $T_{\text{steps}}$.
2:Set model parameter $\xi=v_{0}$.
3:Calculate INAR($\infty$) model parameters based on the scaling limit theory:
4:  $K=1/\zeta(1+\alpha)$
5:  $\delta=K\cdot\Gamma(1-\alpha)/\alpha$
6:  $\mu_{T}=\nu^{*}\cdot\delta^{-1}\cdot T_{\text{steps}}^{\alpha-1}$
7:  $a_{T}=1-(\lambda_{\text{param}}\cdot\delta)/T_{\text{steps}}^{\alpha}$
8:Pre-compute the base kernel $\eta_{n}=K/n^{1+\alpha}$ and its cumulative sum for efficiency.
9:Initialize count array $X^{T}$ of size $T_{\text{steps}}+1$ to zeros.
10:for $n=1$ to $T_{\text{steps}}$ do
11:  Calculate the time-dependent baseline intensity $\hat{\mu}_{T}(n)$ using the formula from Definition 1.
12:  Calculate the self-exciting (convolution) part: $C_{n}=\sum_{s=1}^{n-1}(a_{T}\eta_{n-s})X_{s}^{T}$.
13:  Calculate the total intensity: $\lambda_{n}^{T}=\hat{\mu}_{T}(n)+C_{n}$.
14:  Draw the count for the current step: $X_{n}^{T}\sim\text{Poisson}(\max(\lambda_{n}^{T},0))$.
15:end for
16:Post-processing to obtain the limit process:
17: Compute the cumulative process: $N^{T}_{n}=\sum_{s=1}^{n}X_{s}^{T}$.
18: Apply the scaling factor: $Y_{t}^{T}=\frac{1-a_{T}}{T_{\text{steps}}^{\alpha}\nu^{*}\delta^{-1}}N^{T}_{\lfloor tT_{\text{steps}}\rfloor}$.
19: Compute the numerical derivative to get the final process: $V_{t}=\dot{Y}_{t}^{T}=\frac{\Delta Y^{T}}{\Delta t}$.
20:Output: The simulated path $(V_{t})_{t\in[0,1]}$.
The computational complexity of Algorithm 1 for generating a single path of length $T_{\text{steps}}$ is dominated by the main loop. Inside the loop, the calculation of the convolution term involves a sum of n1n-1 terms, resulting in a total complexity of $\mathcal{O}(T_{\text{steps}}^{2})$ per simulation path. Figure 1 displays a sample path generated using this algorithm, demonstrating its ability to start from the specified initial value.
Figure 1: A sample path of the fractional CIR process $(V_{t}=\dot{Y}_{t})_{0\leq t\leq 1}$, simulated via the INAR($\infty$) approximation. The simulation parameters are $\alpha=0.62$, $\lambda_{\text{param}}=1.0$, $\nu^{*}=200.0$, and a target initial value $v_{0}=0.7$. The number of discrete steps is $T_{\text{steps}}=5000$. The red dashed line indicates the target initial value, which the simulated path closely matches at $t=0$.
3.2 A Goodness-of-Fit Test and application in rough Heston models
Direct validation of the univariate fractional CIR process is challenging due to a scarcity of reliable benchmarks; therefore, we first highlight the framework’s successful application to its well-benchmarked bivariate counterpart, the rough Heston model. To rigorously test the accuracy and goodness-of-fit of our INAR($\infty$) approximation, we focus on the rough volatility case by setting the parameter to $\alpha=0.62$.
The robustness of the underlying INAR($\infty$) framework is highlighted by its successful application in related literature. In the work of Wang and Cui (2025), for instance, a bivariate INAR($\infty$) process is developed to model the interacting dynamics of buy and sell order flows, providing a microstructural foundation for the full rough Heston model. The authors there prove the model’s convergence and validate its accuracy through extensive numerical experiments. The success of the framework in that more complex, bivariate setting (as shown in Wang and Cui (2025)) provides strong collateral support for the validity of our approach. We therefore proceed with a direct numerical validation, comparing our simulation results against the benchmark from Callegaro et al. (2021).
Table 1: Comparison of European option prices for the rough Heston model ($\alpha=0.62$) between the benchmark solution and our INAR($\infty$) simulation. The simulation uses 100,000 paths with $T=250$.
Strike Benchmark Price (Callegaro et al. (2021)) INAR($\infty$) Simulation Price (95% CI)
80 22.1366 22.0152 [21.8835, 22.1469]
90 14.9672 14.8856 [14.7699, 15.0013]
100 9.4737 9.4369 [9.3407, 9.5330]
110 5.6234 5.6300 [5.5542, 5.7058]
120 3.1424 3.1741 [3.1168, 3.2313]
As shown in Table 1, the prices generated by our INAR($\infty$) simulation for the rough case are in excellent agreement with the true prices from the benchmark solution. The differences are statistically insignificant and well within the 95% confidence intervals. This provides strong numerical evidence for the accuracy and practical viability of our approximation method. The specific running code can be found in https://github.com/gagawjbytw/INAR-rough-Heston.
While we use European options in this test for validation against a well-established benchmark, a key advantage of our simulation framework is its versatility. Being a Monte Carlo-based method, it naturally extends to the pricing of path-dependent options—such as Asian, Lookback, and Barrier options—for which analytical or semi-analytical solutions are often intractable. Moreover, the computational efficiency of this approach is a significant benefit. As demonstrated in the bivariate context by Wang and Cui (2025), the INAR($\infty$) simulation framework is inherently parallelizable and proves to be considerably faster than traditional discretization schemes, such as the Euler method, applied to the limiting stochastic Volterra equation. It is stated in Wang and Cui (2025) that the total execution time on a MacBook M2 Pro is 18.3068 seconds for calculating all four options (European, Asian, Lookback, Barrier) for 100000 paths with $T=250$.
4 Proof of Lemmas
4.1 Proof of Lemma 1
Since $\left\lVert\eta\right\rVert_{1}=1$, $\eta$ can be seen as a discrete probability measure. Assume $(G_{i})_{i\geq 1}$ is a sequence of i.i.d. random variables with distribution $\mathbb{P}(G_{i}=n)=\eta_{n}$ for $n\geq 1$. We first consider the rescaled random sum
$J^{T}=\frac{1}{T}\sum_{i=1}^{I^{T}}G_{i},$
where $I^{T}$ is a geometric random variable with parameter $1-a_{T}$. While $J^{T}$ is a discrete random variable, our goal is to show that its distribution converges to a continuous one with a specific density. To this end, we study the Laplace transform of $J^{T}$, denoted by $\hat{\rho}^{T}$, which for $z\geq 0$ satisfies:
$\hat{\rho}^{T}(z)=$ $\mathbb{E}[e^{-zJ^{T}}]=\sum_{k=1}^{\infty}(1-a_{T})a_{T}^{k-1}\mathbb{E}\left[e^{-\frac{z}{T}\sum_{i=1}^{k}G_{i}}\right]$ (22)
$=$ $\sum_{k=1}^{\infty}(1-a_{T})a_{T}^{k-1}\left(\hat{\gamma}\left(\frac{z}{T}\right)\right)^{k}=\frac{(1-a_{T})\hat{\gamma}(z/T)}{1-a_{T}\hat{\gamma}(z/T)},$
where $\hat{\gamma}$ represents the Laplace transform of $G_{1}$, i.e.,
$\hat{\gamma}(z)=\mathbb{E}[e^{-zG_{1}}]=\sum_{n=1}^{\infty}\eta_{n}e^{-zn}.$
By the Abel’s lemma, we obtain
$\sum_{n=0}^{\infty}\eta_{n}e^{-zn}=$ $\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\sum_{k=1}^{n}\eta_{k}=\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\left(1-\sum_{k=n+1}^{\infty}\eta_{k}\right)$
$=$ $1-\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\sum_{k=n+1}^{\infty}\eta_{k}=1-(1-e^{-z})\sum_{n=0}^{\infty}e^{-zn}\sum_{k=n+1}^{\infty}\eta_{k}.$
Since $\eta_{k}\sim\frac{K}{k^{\alpha+1}}$ for sufficiently large $k$, it follows that $\sum_{k=n+1}^{\infty}\eta_{k}\sim\frac{K}{\alpha}n^{-\alpha}$. From Lemma 3, we obtain
$\sum_{n=0}^{\infty}\left(e^{-zn}-e^{-z(n+1)}\right)\sum_{k=n+1}^{\infty}\eta_{k}\sim\frac{K}{\alpha}\Gamma(1-\alpha)z^{\alpha},\quad\text{as }z\downarrow 0.$
This yields the asymptotic expansion for $\hat{\gamma}(z)$ around the origin:
$\hat{\gamma}(z)=1-K\frac{\Gamma(1-\alpha)}{\alpha}z^{\alpha}+o(z^{\alpha}),\quad\text{as }z\downarrow 0.$
Let us define
$v_{T}=\delta^{-1}T^{\alpha}(1-a_{T}),\quad\text{where}\quad\delta=K\frac{\Gamma(1-\alpha)}{\alpha}.$ (23)
Under Assumption 2, this sequence converges to a limit, which we denote by $v_{\infty}$:
$\lim_{T\to\infty}v_{T}=\lim_{T\to\infty}\delta^{-1}T^{\alpha}(1-a_{T})=\delta^{-1}(\lambda\delta)=\lambda=:v_{\infty}.$
Substituting the expansion of $\hat{\gamma}(z/T)$ into the expression (22) for $\hat{\rho}^{T}(z)$ and taking the limit as $T\to\infty$, we find that for any fixed $z\geq 0$:
$\lim_{T\to\infty}\hat{\rho}^{T}(z)=\frac{\lambda}{\lambda+z^{\alpha}}=\frac{v_{\infty}}{v_{\infty}+z^{\alpha}}.$
The right-hand side is the Laplace transform of the probability density function $x\mapsto\lambda x^{\alpha-1}E_{\alpha,\alpha}(-\lambda x^{\alpha})$. By Lévy’s continuity theorem for Laplace transforms, the convergence of the transforms implies the weak convergence of the corresponding probability measures. This completes the proof of the lemma.
4.2 Proof of Lemma 2
The proof consists of two main steps. First, we rigorously show that for a function $f(t):=t^{\alpha}l(t)$, which is regularly varying with index $\alpha>-1$, its sum is asymptotically equivalent to its integral. Second, we apply Karamata’s Integral Theorem.
First, let’s prove that $\sum_{t=X}^{x}f(t)\sim\int_{X}^{x}f(t)dt$. Our goal is to prove that $\lim_{x\to\infty}\left(\sum_{t=X}^{x}f(t)\right)/\left(\int_{X}^{x}f(t)dt\right)=1$. We use the Stolz–Cesàro theorem. Let $a_{x}=\sum_{t=X}^{x}f(t)$ and $b_{x}=\int_{X}^{x}f(t)dt$. Since $\alpha>-1$, the function $f(t)$ is eventually positive, and thus $b_{x}$ is strictly increasing to infinity, satisfying the conditions for the theorem. We therefore analyze the limit of the ratio of differences:
$\lim_{x\to\infty}\frac{a_{x}-a_{x-1}}{b_{x}-b_{x-1}}=\lim_{x\to\infty}\frac{f(x)}{\int_{x-1}^{x}f(t)dt}.$
To evaluate this limit, we show that the denominator is asymptotically equivalent to the numerator. For any $t\in[x-1,x]$, we can write $t=\lambda x$, where $\lambda\in[1-1/x,1]$. As $x\to\infty$, this interval for $\lambda$ converges to the point $\{1\}$.
According to the Uniform Convergence Theorem for regularly varying functions (Bingham et al., 1989, Theorem 1.5.2), the convergence of $f(\lambda x)/f(x)\to\lambda^{\alpha}$ is uniform for $\lambda$ in any compact interval $[a,b]$ with $0<a\leq b<\infty$. Let us choose such an interval that contains $1$, for example $[1/2,2]$. For $x$ large enough, our interval [11/x,1][1-1/x,1] is contained within $[1/2,2]$.
The uniform convergence implies that for any $\varepsilon>0$, there exists an $N$ such that for all $x>N$ and for all $t\in[x-1,x]$ (and thus for all corresponding $\lambda\in[1-1/x,1]$), we have:
$(1-\varepsilon)f(x)\leq f(t)\leq(1+\varepsilon)f(x).$
Integrating this inequality over the interval [x1,x][x-1,x], which has length 1, yields:
$(1-\varepsilon)f(x)\leq\int_{x-1}^{x}f(t)dt\leq(1+\varepsilon)f(x).$
This demonstrates that $\int_{x-1}^{x}f(t)dt\sim f(x)$ as $x\to\infty$. Therefore, the limit of the ratio of differences is 1. By the Stolz-Cesàro theorem, the limit of the original ratio is also 1, which proves that $\sum_{t=X}^{x}f(t)\sim\int_{X}^{x}f(t)dt$.
Next, we apply Karamata’s Integral Theorem. Furthermore, Karamata’s theorem for integrals of regularly varying functions (Bingham et al., 1989, Proposition 1.5.8) states that
$\int_{X}^{x}t^{\alpha}l(t)dt\sim\frac{1}{\alpha+1}x^{\alpha+1}l(x)\quad\text{as }x\to\infty.$
Combining the above asymptotic equivalences, we obtain the desired result:
$\sum_{t=X}^{x}t^{\alpha}l(t)\sim\int_{X}^{x}t^{\alpha}l(t)dt\sim\frac{1}{\alpha+1}x^{\alpha+1}l(x).$
This completes the proof.
4.3 Proof of Lemma 3
Our goal is to connect the asymptotic behavior of the sequence $(\beta_{n})$ to its transform $\hat{\beta}(s)$. We achieve this via the sequence of partial sums, $V_{n}=\sum_{k=0}^{n-1}\beta_{k}$ (with $V_{0}=0$).
First, let us establish the exact relationship between the transform of the sequence, $\hat{\beta}(s)$, and the transform of its partial sums, $\hat{V}(s)$. Following the definitions, we have:
$\hat{V}(s)$ $=\sum_{n=0}^{\infty}(e^{-sn}-e^{-s(n+1)})\sum_{m=0}^{n-1}\beta_{m}$
$=\sum_{m=0}^{\infty}\beta_{m}\sum_{n=m+1}^{\infty}(e^{-sn}-e^{-s(n+1)})$
$=\sum_{m=0}^{\infty}\beta_{m}e^{-s(m+1)}$
$=\frac{e^{-s}}{1-e^{-s}}\hat{\beta}(s).$
As $s\downarrow 0$, we have $e^{-s}\to 1$ and $1-e^{-s}\sim s$. This implies the asymptotic relationship $\hat{V}(s)\sim s^{-1}\hat{\beta}(s)$.
Now, to find the asymptotic behavior of $\hat{V}(s)$, we first need the asymptotics of $V_{n}$. We are given that $\beta_{k}\sim\frac{ck^{\rho}l(k)}{\Gamma(1+\rho)}$. This is precisely the form required by Lemma 2. Applying the lemma, we find the asymptotic behavior of the partial sums:
$V_{n}=\sum_{k=0}^{n-1}\beta_{k}\sim\frac{c}{\Gamma(1+\rho)}\frac{n^{\rho+1}l(n)}{\rho+1}=\frac{cn^{\rho+1}l(n)}{\Gamma(\rho+2)}.$
With the behavior of $V_{n}$ established, we can apply Karamata’s Tauberian Theorem (e.g., Corollary 1.7.3 in Bingham et al. (1989)) to connect the asymptotics of the partial sums to their transform. The theorem yields:
$\hat{V}(s)\sim cs^{-(\rho+1)}l(1/s)\quad\text{as }s\downarrow 0.$
Finally, combining this with our derived relationship $\hat{\beta}(s)\sim s\hat{V}(s)$, we obtain the desired result:
$\hat{\beta}(s)\sim cs^{-\rho}l(1/s)\quad\text{as }s\downarrow 0.$
This completes the proof.
4.4 Proof of Lemma 4
We use the criterion from Lemma 5 to verify the $C$-tightness for the sequence of càdlàg processes $(Y_{t}^{T})_{t\in[0,1]}$.
Before verifying the conditions, we establish the asymptotic orders of the key moments for large $T$ and for any time index $s$ such that $s/T\to t\in(0,1]$. We use the decomposition $\lambda_{s}^{T}=q_{s}^{T}+\kappa_{s}^{T}$, where $q_{s}^{T}$ is the deterministic mean and $\kappa_{s}^{T}$ is the zero-mean component.
  • We first calculate the first Moment of $\lambda_{s}^{T}$. We have
    $\mathbb{E}[\lambda_{s}^{T}]=q_{s}^{T}=\mu^{T}+\mu^{T}\sum_{u=1}^{s-1}A_{u}^{T}\leq\frac{\mu^{T}}{1-a_{T}}\sim\frac{\nu^{*}}{(\lambda\delta^{2})}T^{2\alpha-1}.$
    Hence, there exists a constant $\bar{C}_{1}>0$ such that $\mathbb{E}[\lambda_{s}^{T}]\leq\bar{C}_{1}T^{2\alpha-1}$ for $T\geq 1$.
  • Next, we calculate the second Moment of $\lambda_{s}^{T}$. The second moment is $\mathbb{E}[(\lambda_{s}^{T})^{2}]=(q_{s}^{T})^{2}+\mathbb{E}[(\kappa_{s}^{T})^{2}]$. We now rigorously derive the order of $\mathbb{E}[(\kappa_{s}^{T})^{2}]$. The exact expression is:
    $\mathbb{E}[(\kappa_{s}^{T})^{2}]=\sum_{k=1}^{s-1}(A_{k}^{T})^{2}\mathbb{E}[\lambda_{s-k}^{T}].$
    Our goal is to prove that $\mathbb{E}[(\kappa_{s}^{T})^{2}]\leq CT^{4\alpha-2}$ for some constant $C>0$ as $T\to\infty$ and $s/T\to t\in(0,1]$.
    First, we calculate the asymptotic order of $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$. From the proof of Lemma 1, we established that the scaled sequence $TA_{k}^{T}/\left\lVert A^{T}\right\rVert_{1}$ converges to a limiting density. For large $k$ and $T$ such that $k/T\to x>0$, this implies:
    $A_{k}^{T}\sim\frac{\left\lVert A^{T}\right\rVert_{1}}{T}\left(\lambda(k/T)^{\alpha-1}E_{\alpha,\alpha}(-\lambda(k/T)^{\alpha})\right).$
    From Assumption 2, we have the following asymptotics:
    $\left\lVert A^{T}\right\rVert_{1}=\frac{a_{T}}{1-a_{T}}\sim\frac{1}{1-a_{T}}\sim\frac{T^{\alpha}}{\lambda\delta}.$
    Substituting this into the expression for $A_{k}^{T}$:
    $A_{k}^{T}$ $\sim\frac{T^{\alpha}}{\lambda\delta T}\left(\lambda(k/T)^{\alpha-1}E_{\alpha,\alpha}(-\lambda(k/T)^{\alpha})\right)$
    $=\frac{T^{\alpha-1}}{\delta}\cdot T^{1-\alpha}k^{\alpha-1}E_{\alpha,\alpha}(-\lambda(k/T)^{\alpha})$
    $=\frac{1}{\delta}k^{\alpha-1}E_{\alpha,\alpha}(-\lambda(k/T)^{\alpha}).$
    For large $k$, the Mittag-Leffler function term is bounded. This shows that $A_{k}^{T}$ is regularly varying with index $\alpha-1$. Consequently, $(A_{k}^{T})^{2}$ is regularly varying with index $2\alpha-2$.
    To determine the asymptotic order of the sum more rigorously, we first approximate it by its corresponding integral. Using the asymptotic form of $A_{k}^{T}$ derived from the proof of Lemma 1, for $s=tT$, we have:
    $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$ $\sim\int_{1}^{s-1}(A_{x}^{T})^{2}dx$
    $\approx\int_{1}^{tT}\left(\frac{1}{\delta}x^{\alpha-1}E_{\alpha,\alpha}(-\lambda(x/T)^{\alpha})\right)^{2}dx$
    $=\frac{1}{\delta^{2}}\int_{1}^{tT}x^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda(x/T)^{\alpha})\right]^{2}dx.$
    To properly evaluate the dependence on $T$, we perform a change of variables by letting $u=x/T$. This implies $x=uT$ and dx=Tdudx=T\,du. The limits of integration transform from $[1,tT]$ to $[1/T,t]$. Substituting these into the integral yields:
    $\sum_{k=1}^{s-1}(A_{k}^{T})^{2}$ $\approx\frac{1}{\delta^{2}}\int_{1/T}^{t}(uT)^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda u^{\alpha})\right]^{2}(T\,du)$
    $=\frac{T^{2\alpha-1}}{\delta^{2}}\int_{1/T}^{t}u^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda u^{\alpha})\right]^{2}du.$
    As $T\to\infty$, the lower limit of the integral $1/T\to 0$. The integral itself converges to a finite, positive constant that depends on $t$ but not on $T$:
    $\lim_{T\to\infty}\int_{1/T}^{t}u^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda u^{\alpha})\right]^{2}du=\int_{0}^{t}u^{2\alpha-2}\left[E_{\alpha,\alpha}(-\lambda u^{\alpha})\right]^{2}du=:C_{*}(t).$
    This limit is well-defined and finite because as $u\to 0$, the integrand behaves like $u^{2\alpha-2}$, and since $\alpha>1/2$, the exponent $2\alpha-2>-1$, ensuring the integral converges at the origin. Thus, for any $t\in[0,1]$, we have established the asymptotic order:
    $\sum_{k=1}^{\lfloor tT\rfloor-1}(A_{k}^{T})^{2}\sim\frac{C_{*}(t)}{\delta^{2}}T^{2\alpha-1},$
    where
    $C_{*}(t)=\int_{0}^{t}u^{2\alpha-2}[E_{\alpha,\alpha}(-\lambda u^{\alpha})]^{2}du.$
    Since the integrand is non-negative, $C_{*}(t)$ is a non-decreasing function of $t$. This implies that for any $t\in[0,1]$, $C_{*}(t)\leq C_{*}(1)$. Therefore, there exists a uniform constant $\bar{C}_{A}:=C_{*}(1)/\delta^{2}$, independent of $t$ and $T$, such that for all sufficiently large $T$:
    $\sup_{t\in[0,1]}\left(T^{-(2\alpha-1)}\sum_{k=1}^{\lfloor tT\rfloor-1}(A_{k}^{T})^{2}\right)\leq\bar{C}_{A}.$
    This provides the necessary uniform bound for the sum.
    Next, we establish a non-asymptotic upper bound for $\mathbb{E}[(\kappa_{s}^{T})^{2}]$.
    First, from our preliminary calculations, for all $T\geq 1$ and $j\geq 1$, there exists a uniform constant $\bar{C}_{1}$ such that $\mathbb{E}[\lambda_{j}^{T}]\leq\bar{C}_{1}T^{2\alpha-1}$.
    Furthermore, the analysis in the previous paragraph establishes that there exists a constant $\bar{C}_{2}$, independent of both $t$ and $T$, such that for any $t\in[0,1]$ and all $T\geq 1$:
    $\sum_{k=1}^{\lfloor tT\rfloor-1}(A_{k}^{T})^{2}\leq\bar{C}_{2}T^{2\alpha-1}.$
    (This uniform constant $\bar{C}_{2}$ is obtained by taking the bound corresponding to $t=1$, since the underlying integral defining the bound is non-decreasing in $t$).
    Applying these uniform bounds for $s=\lfloor tT\rfloor$, we get:
    $\mathbb{E}[(\kappa_{s}^{T})^{2}]$ $=\sum_{k=1}^{s-1}(A_{k}^{T})^{2}\mathbb{E}[\lambda_{s-k}^{T}]$
    $\leq\left(\sup_{j\geq 1}\mathbb{E}[\lambda_{j}^{T}]\right)\left(\sum_{k=1}^{s-1}(A_{k}^{T})^{2}\right)$
    $\leq(\bar{C}_{1}T^{2\alpha-1})\cdot(\bar{C}_{2}T^{2\alpha-1})$
    $=\bar{C}_{1}\bar{C}_{2}T^{4\alpha-2}.$
    This provides the required non-asymptotic upper bound for the second moment.
    Since $(q_{s}^{T})^{2}=(\mathbb{E}[\lambda_{s}^{T}])^{2}\leq(\bar{C}_{1})^{2}T^{4\alpha-2}$, we finally conclude that
    $\mathbb{E}[(\lambda_{s}^{T})^{2}]\leq\left((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)\cdot T^{4\alpha-2}.$
  • Second Moment of $X_{s}^{T}$: From $\mathbb{E}[(X_{s}^{T})^{2}]=\mathbb{E}[\lambda_{s}^{T}]+\mathbb{E}[(\lambda_{s}^{T})^{2}]$, we have
    $\mathbb{E}[(X_{s}^{T})^{2}]\leq\bar{C}_{1}T^{2\alpha-1}+\left((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}\leq\left(\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}$ (24)
    for $T\geq 1$.
Now let us verify the conditions for $C$-tightness.
  1. 1.
    Condition 1 (Initial Value): Since $Y_{0}^{T}=0$, this condition is trivially satisfied.
  2. 2.
    Condition 2 (Small Increments in Probability): Given $\theta>2$, for large $T$, the sum $\sum_{s=\lfloor kT^{1-\theta}\rfloor+1}^{\lfloor(k+1)T^{1-\theta}\rfloor}X_{s}^{T}$ contains at most one term. This justifies:
    $\sup_{k,h}\left|\Delta_{h}Y_{k/T^{\theta}}^{T}\right|\leq\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sup_{s\in\{1,\dots,\lfloor T\rfloor\}}X_{s}^{T}.$
    Let $C_{T}(\epsilon)=\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}\epsilon=\mathcal{O}(T^{2\alpha})$. There exists a constant $\bar{C}_{3}>0$ such that $\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}\geq\bar{C}_{3}T^{2\alpha}$ for $T\geq 1$. By the union bound and Markov’s inequality:
    $\mathbb{P}\left(\sup_{s\leq\lfloor T\rfloor}X_{s}^{T}>C_{T}(\epsilon)\right)\leq$ $\sum_{s=0}^{\lfloor T\rfloor}\frac{\mathbb{E}[(X_{s}^{T})^{2}]}{C_{T}(\epsilon)^{2}}$
    $\leq$ $T\cdot\frac{\left(\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}}{\bar{C}_{3}^{2}T^{4\alpha}}$
    $=$ $\frac{\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}}{\bar{C}_{3}^{2}}T^{-1}\to 0.$
    Thus, Condition 2 is satisfied.
  3. 3.
    Condition 3 (Moment Bound on Increments): First, there exists a constant $\bar{C}_{5}>0$ such that for any $T\geq 1$ and $t,h\in[0,1]$:
    $(\lfloor T(t+h)\rfloor-\lfloor Tt\rfloor+1)^{2}\leq\bar{C}_{5}T^{2}h^{2}.$
    Then by the Cauchy-Schwarz inequality:
    $\mathbb{E}\left[\left|\Delta_{h}Y_{t}^{T}\right|^{2}\right]\leq$ $\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}(\lfloor T(t+h)\rfloor-\lfloor Tt\rfloor+1)\sum_{i=\lfloor Tt\rfloor}^{\lfloor T(t+h)\rfloor}\mathbb{E}\left[(X_{i}^{T})^{2}\right]$
    $\leq$ $\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}(\lfloor T(t+h)\rfloor-\lfloor Tt\rfloor+1)^{2}\max_{i=\lfloor Tt\rfloor,\dots,\lfloor T(t+h)\rfloor}\mathbb{E}\left[(X_{i}^{T})^{2}\right].$
    From (24), $\mathbb{E}[(X_{i}^{T})^{2}]\leq\left(\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}\right)T^{4\alpha-2}$. Hence,
    $\mathbb{E}\left[\left|\Delta_{h}Y_{t}^{T}\right|^{2}\right]\leq\frac{\bar{C}_{5}(\bar{C}_{1}+(\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2})}{\bar{C}_{3}^{2}}h^{2}.$
    This demonstrates the condition holds with $p=2,b_{1}=2,a_{1}=0$. Thus, Condition 3 is satisfied.
4.5 Proof of Lemma 6
We aim to show that for any $\epsilon>0$, $\mathbb{P}(\sup_{t\in[0,1]}|Y_{t}^{T}-\Lambda_{t}^{T}|>\epsilon)\to 0$ as $T\to\infty$. We have
$Y_{t}^{T}-\Lambda_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}M_{\lfloor tT\rfloor}^{T},$
where $M_{n}^{T}=\sum_{s=1}^{n}(X_{s}^{T}-\lambda_{s}^{T})$ is a zero-mean martingale.
Applying Doob’s $L^{2}$-inequality to the martingale $M^{T}$, we get:
$\mathbb{E}\left[\sup_{t\in[0,1]}\left(Y_{t}^{T}-\Lambda_{t}^{T}\right)^{2}\right]\leq 4\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}\mathbb{E}\left[(M_{\lfloor T\rfloor}^{T})^{2}\right]\$
The second moment of the martingale $M_{\lfloor T\rfloor}^{T}$ is equal to the expectation of its predictable quadratic variation:
$\mathbb{E}\left[(M_{\lfloor T\rfloor}^{T})^{2}\right]$ $=\mathbb{E}\left[\left(\sum_{s=1}^{\lfloor T\rfloor}(X_{s}^{T}-\lambda_{s}^{T})\right)^{2}\right]$
$=\mathbb{E}\left[\sum_{s=1}^{\lfloor T\rfloor}(X_{s}^{T}-\lambda_{s}^{T})^{2}\right]$ (Cross-terms vanish for martingale differences)
$=\sum_{s=1}^{\lfloor T\rfloor}\mathbb{E}\left[\mathbb{E}[(X_{s}^{T}-\lambda_{s}^{T})^{2}|\mathcal{F}_{s-1}]\right]$
$=\sum_{s=1}^{\lfloor T\rfloor}\mathbb{E}[\lambda_{s}^{T}]$ $\text{(Since }X_{s}^{T}|\mathcal{F}_{s-1}\sim\text{Poisson}(\lambda_{s}^{T})\text{)}$
$=\mathbb{E}\left[\sum_{s=1}^{\lfloor T\rfloor}X_{s}^{T}\right]=\mathbb{E}[N_{\lfloor T\rfloor}^{T}].$ $\text{(Since }\mathbb{E}[X_{s}^{T}]=\mathbb{E}[\lambda_{s}^{T}]\text{)}$
We now establish the asymptotic behavior of $\mathbb{E}[N_{\lfloor T\rfloor}^{T}]$. The expectation $\mathbb{E}[X_{s}^{T}]$ satisfies the discrete renewal equation $\mathbb{E}[X_{s}^{T}]=\mu^{T}+\sum_{k=1}^{s-1}\eta_{k}^{T}\mathbb{E}[X_{s-k}^{T}]$. Since $\|\eta^{T}\|_{1}=a_{T}<1$, as $s\to\infty$, $\mathbb{E}[X_{s}^{T}]$ converges to the stationary mean $\mu^{T}/(1-a_{T})$. By Cesàro’s theorem on the mean of convergent sequences, we have:
$\frac{1}{n}\mathbb{E}[N_{n}^{T}]=\frac{1}{n}\sum_{s=1}^{n}\mathbb{E}[X_{s}^{T}]\to\frac{\mu^{T}}{1-a_{T}}\quad\text{as }n\to\infty.$
Thus, for $n=\lfloor T\rfloor$, we have the asymptotic equivalence $\mathbb{E}[N_{\lfloor T\rfloor}^{T}]\sim\frac{\mu^{T}T}{1-a_{T}}$. Substituting this back, we get:
$\mathbb{E}\left[\sup_{t\in[0,1]}\left(Y_{t}^{T}-\Lambda_{t}^{T}\right)^{2}\right]$ $\leq C\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}\left(\frac{\mu^{T}T}{1-a_{T}}\right)$
$\leq C^{\prime}T^{-2\alpha}$
for some constants $C$ and $C^{\prime}$. So this term converges to 0 as $T\to\infty$. Convergence in $L^{2}$ implies convergence in probability, which completes the proof.
4.6 Proof of Lemma 7
We consider the generating functions (or $z$-transforms) of the sequences involved. Since $(\eta_{n})_{n\geq 1}$ is a non-negative sequence in $\ell^{1}$, its generating function $\mathcal{G}(\eta)(z)=\sum_{n=1}^{\infty}\eta_{n}z^{n}$ converges absolutely for all complex $z$ with $|z|\leq 1$. All subsequent manipulations are valid for $|z|<1$, where $|\mathcal{G}(\eta)(z)|<\sum_{n=1}^{\infty}\eta_{n}\leq 1$.
Taking the $z$-transform on both sides of equation (11) and using the convolution property Oppenheim (1999), we obtain:
$\mathcal{G}(x)(z)=\mathcal{G}(y)(z)+\mathcal{G}(\eta)(z)\cdot\mathcal{G}(x)(z).$
Solving for $\mathcal{G}(x)(z)$, we get:
$\mathcal{G}(x)(z)=\frac{\mathcal{G}(y)(z)}{1-\mathcal{G}(\eta)(z)}=\mathcal{G}(y)(z)\cdot\left(1+\frac{\mathcal{G}(\eta)(z)}{1-\mathcal{G}(\eta)(z)}\right).$
To interpret the second term, we find the generating function of the sequence $(A_{n})_{n\geq 1}$. By definition, $A_{n}=\sum_{k=1}^{\infty}(\eta^{*k})_{n}$, where $\eta^{*k}$ is the $k$-fold convolution of $\eta$. The generating function of a convolution is the product of the generating functions. Thus, the generating function of $\eta^{*k}$ is $(\mathcal{G}(\eta)(z))^{k}$. The generating function of $A_{n}$ is therefore:
$\mathcal{G}(A)(z)$ $=\sum_{n=1}^{\infty}A_{n}z^{n}=\sum_{n=1}^{\infty}\left(\sum_{k=1}^{\infty}(\eta^{*k})_{n}\right)z^{n}$
$=\sum_{k=1}^{\infty}\sum_{n=1}^{\infty}(\eta^{*k})_{n}z^{n}=\sum_{k=1}^{\infty}(\mathcal{G}(\eta)(z))^{k}.$
This is a geometric series, which for $|\mathcal{G}(\eta)(z)|<1$ converges to:
$\mathcal{G}(A)(z)=\frac{\mathcal{G}(\eta)(z)}{1-\mathcal{G}(\eta)(z)}.$
Substituting this back into the expression for $\mathcal{G}(x)(z)$, we have:
$\mathcal{G}(x)(z)=\mathcal{G}(y)(z)\cdot(1+\mathcal{G}(A)(z))=\mathcal{G}(y)(z)+\mathcal{G}(y)(z)\mathcal{G}(A)(z).$
By performing the inverse $z$-transform and applying the convolution theorem again, we obtain the unique solution:
$x_{n}=y_{n}+\sum_{i=1}^{n-1}A_{i}y_{n-i}.$
This concludes the proof.
4.7 Proof of Lemma 8
We notice that
$\lambda_{n}^{T}=$ $\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}X_{s}^{T}=\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}+\lambda_{s}^{T}\right)$
$=$ $\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right)+\sum_{s=1}^{n-1}\eta_{n-s}^{T}\lambda_{s}^{T}.$
By the Lemma 7, we obtain
$\lambda_{n}^{T}=$ $\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right)+\sum_{s=1}^{n-1}A_{n-s}^{T}\left(\mu^{T}+\sum_{u=1}^{s}\eta_{s-u}^{T}\left(M_{u}^{T}-M_{u-1}^{T}\right)\right)$
$=$ $\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right)+\mu^{T}\sum_{s=1}^{n-1}A_{n-s}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}\left(\sum_{u=1}^{s-1}\eta_{s-u}^{T}\left(M_{u}^{T}-M_{u-1}^{T}\right)\right)$
$=$ $\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right)+\mu^{T}\sum_{s=1}^{n-1}A_{n-s}^{T}+\sum_{u=1}^{n-1}(M_{u}^{T}-M_{u-1}^{T})\sum_{s=u+1}^{n-1}\eta_{s-u}^{T}A_{n-s}^{T}$
$=$ $\mu^{T}+\sum_{s=1}^{n-1}\eta_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right)+\mu^{T}\sum_{s=1}^{n-1}A_{n-s}^{T}+\sum_{u=1}^{n-1}(M_{u}^{T}-M_{u-1}^{T})\left(A_{n-u}^{T}-\eta_{n-u}^{T}\right)$
$=$ $\mu^{T}+\mu^{T}\sum_{s=1}^{n-1}A_{n-s}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right).$
4.8 Proof of Lemma 9
The proof relies on the application of the discrete renewal equation from Lemma 7.
First, let’s prove Equation (19). By taking the expectation and using the law of total expectation:
$\mathbb{E}[N_{n}^{T}]$ $=\sum_{i=1}^{n}\mathbb{E}[X_{i}^{T}]=\sum_{i=1}^{n}\mathbb{E}[\lambda_{i}^{T}]$
$=\sum_{i=1}^{n}\mathbb{E}\left[\hat{\mu}_{T}(i)+\sum_{s=1}^{i-1}\eta_{s}^{T}X_{i-s}^{T}\right]$
$=\sum_{i=1}^{n}\hat{\mu}_{T}(i)+\sum_{i=1}^{n}\sum_{s=1}^{i-1}\eta_{s}^{T}\mathbb{E}[X_{i-s}^{T}]$
$=\sum_{i=1}^{n}\hat{\mu}_{T}(i)+\sum_{s=1}^{n-1}\eta_{s}^{T}\sum_{i=s+1}^{n}\mathbb{E}[X_{i-s}^{T}]$
$=\sum_{i=1}^{n}\hat{\mu}_{T}(i)+\sum_{s=1}^{n-1}\eta_{s}^{T}\mathbb{E}[N_{n-s}^{T}].$
This is a discrete renewal equation for $x_{n}=\mathbb{E}[N_{n}^{T}]$ with driving term $y_{n}=\sum_{i=1}^{n}\hat{\mu}_{T}(i)$. Applying the solution from Lemma 7, we have:
$\mathbb{E}[N_{n}^{T}]=\sum_{i=1}^{n}\hat{\mu}_{T}(i)+\sum_{s=1}^{n-1}A_{s}^{T}\left(\sum_{i=1}^{n-s}\hat{\mu}_{T}(i)\right).$
Re-arranging the order of summation in the second term gives the desired result.
Next, for Equation (20), we start by expressing $N_{n}^{T}$ in terms of its martingale and intensity components:
$N_{n}^{T}=\sum_{i=1}^{n}X_{i}^{T}=\sum_{i=1}^{n}(M_{i}^{T}-M_{i-1}^{T}+\lambda_{i}^{T})=M_{n}^{T}+\sum_{i=1}^{n}\lambda_{i}^{T}.$
Subtracting the expectation, we get:
$N_{n}^{T}-\mathbb{E}[N_{n}^{T}]=M_{n}^{T}+\sum_{i=1}^{n}(\lambda_{i}^{T}-\mathbb{E}[\lambda_{i}^{T}]).$
Let $Z_{n}=N_{n}^{T}-\mathbb{E}[N_{n}^{T}]$. Then $\lambda_{i}^{T}-\mathbb{E}[\lambda_{i}^{T}]=\sum_{s=1}^{i-1}\eta_{s}^{T}(X_{i-s}^{T}-\mathbb{E}[X_{i-s}^{T}])=\sum_{s=1}^{i-1}\eta_{s}^{T}Z_{i-s}$. The equation becomes a discrete renewal equation for $Z_{n}$:
$Z_{n}=M_{n}^{T}+\sum_{i=1}^{n}\sum_{s=1}^{i-1}\eta_{s}^{T}Z_{i-s}=M_{n}^{T}+\sum_{s=1}^{n-1}\eta_{s}^{T}Z_{n-s}.$
Applying the solution from Lemma 7 with driving term $y_{n}=M_{n}^{T}$ gives the result:
$N_{n}^{T}-\mathbb{E}[N_{n}^{T}]=M_{n}^{T}+\sum_{s=1}^{n-1}A_{s}^{T}M_{n-s}^{T}=M_{n}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}M_{s}^{T}.$
This completes the proof.
5 Proof of the Main Results
5.1 Proof of Proposition 1
The proof proceeds in three main steps. First, we establish the joint tightness of the sequence $(Z^{T},Y^{T})$. Second, we show that the quadratic variation $Z^{T},Z^{T}]_{t}$ converges to the same limit as $Y^{T}_{t}$. Finally, we use this to characterize the limit process $(Z,Y)$.
1. Tightness of $(Z^{T},Y^{T})$. From Lemma 4, we have already established that the sequences $(Y^{T})$ and $(\Lambda^{T})$ are $\mathbb{C}$-tight. The process $Z^{T}$ is defined as
$Z_{t}^{T}=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}(Y_{t}^{T}-\Lambda_{t}^{T}).$
Since a linear combination of $\mathbb{C}$-tight sequences is also $\mathbb{C}$-tight, $(Z^{T})$ is $\mathbb{C}$-tight. By Theorem VI-4.13 in Jacod and Shiryaev (2013), this implies that the sequence $(Z^{T})$ is tight in the Skorokhod space. As $(Y^{T})$ is also tight (from $\mathbb{C}$-tightness), the sequence of pairs $(Z^{T},Y^{T})$ is jointly tight.
2. Convergence of the Quadratic Variation. Let $(Z,Y)$ be any limit point of $(Z^{T},Y^{T})$. To prove that $Y=[Z,Z]$, we follow the standard procedure of showing that both $Z^{T},Z^{T}]_{t}$ and $Y^{T}_{t}$ converge in probability to the same process. We do this by showing that
$Z^{T},Z^{T}]_{t}-\langle Z^{T},Z^{T}\rangle_{t}\to 0$
and
$\langle Z^{T},Z^{T}\rangle_{t}-Y^{T}_{t}\to 0.$
Let $C_{T}^{2}=\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}$ and $D_{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}$. Note that
$Z_{t}^{T}=C_{T}(Y_{t}^{T}-\Lambda_{t}^{T})=C_{T}D_{T}M_{\lfloor tT\rfloor}^{T}.$
The increment is $\Delta Z_{s/T}^{T}=C_{T}D_{T}(X_{s}^{T}-\lambda_{s}^{T})$.
(a) Convergence of $Z^{T},Z^{T}]_{t}$ to $\langle Z^{T},Z^{T}\rangle_{t}$. The predictable quadratic variation of $Z^{T}$ is
$\langle Z^{T},Z^{T}\rangle_{t}=$ $\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}[(\Delta Z_{s/T}^{T})^{2}|\mathcal{F}_{s-1}]$
$=$ $(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}[(X_{s}^{T}-\lambda_{s}^{T})^{2}|\mathcal{F}_{s-1}]$
$=$ $(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}.$
The sharp bracket (quadratic variation) is
$Z^{T},Z^{T}]_{t}=\sum_{s=1}^{\lfloor tT\rfloor}(\Delta Z_{s/T}^{T})^{2}=(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}(X_{s}^{T}-\lambda_{s}^{T})^{2}.$
Consider the process
$U_{n}^{T}:=[Z^{T},Z^{T}]_{n/T}-\langle Z^{T},Z^{T}\rangle_{n/T}.$
This is a martingale. We show it converges to zero in $L^{2}$:
$\mathbb{E}[(U_{\lfloor tT\rfloor}^{T})^{2}]$ $=(C_{T}D_{T})^{4}\mathbb{E}\left[\left(\sum_{s=1}^{\lfloor tT\rfloor}\left((X_{s}^{T}-\lambda_{s}^{T})^{2}-\lambda_{s}^{T}\right)\right)^{2}\right]$
$=(C_{T}D_{T})^{4}\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}\left[\left((X_{s}^{T}-\lambda_{s}^{T})^{2}-\lambda_{s}^{T}\right)^{2}\right]$
$=(C_{T}D_{T})^{4}\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}[\text{Var}((X_{s}^{T}-\lambda_{s}^{T})^{2}|\mathcal{F}_{s-1})]$
$=(C_{T}D_{T})^{4}\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}[\lambda_{s}^{T}+2(\lambda_{s}^{T})^{2}].$
To show this convergence, we analyze the expectation of the sum. From the moment calculations in the proof of Lemma 4, we know there exist uniform constants $\bar{C}_{1},\bar{C}_{2}$ such that for all $s$ and $T$:
$\mathbb{E}[\lambda_{s}^{T}]\leq\bar{C}_{1}T^{2\alpha-1}\quad\text{and}\quad\mathbb{E}[(\lambda_{s}^{T})^{2}]\leq((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2})T^{4\alpha-2}.$
Therefore, the sum can be bounded as follows:
$\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}[\lambda_{s}^{T}+2(\lambda_{s}^{T})^{2}]$ $\leq\sum_{s=1}^{\lfloor tT\rfloor}(\bar{C}_{1}T^{2\alpha-1}+2((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2})T^{4\alpha-2})$
$\leq\lfloor tT\rfloor\cdot(\bar{C}_{1}T^{2\alpha-1}+2((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2})T^{4\alpha-2})$
$\leq(\bar{C}_{1}+2((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}))T^{4\alpha-1}\quad(\text{since }2\alpha-1<4\alpha-2\ \text{and}\ t\in[0,1]).$
For the pre-factor, by Assumption 2, there exists a constant $\bar{C}_{6}$ such that for all sufficiently large $T$:
$(C_{T}D_{T})^{4}=D_{T}^{2}=\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}\leq\bar{C}_{6}T^{-4\alpha}.$
Combining these non-asymptotic bounds, we have:
$\mathbb{E}[(U_{\lfloor tT\rfloor}^{T})^{2}]$ $\leq(\bar{C}_{6}T^{-4\alpha})\cdot\left((\bar{C}_{1}+2((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}))T^{4\alpha-1}\right)$
$=\bar{C}_{6}(\bar{C}_{1}+2((\bar{C}_{1})^{2}+\bar{C}_{1}\bar{C}_{2}))T^{-1}.$
As $T\to\infty$, this upper bound for the $L^{2}$-norm converges to 0. Convergence in $L^{2}$ implies convergence in probability, thus $Z^{T},Z^{T}]_{t}-\langle Z^{T},Z^{T}\rangle_{t}\to 0$ in probability, uniformly in $t$ by Doob’s inequality.
(b) Convergence of $\langle Z^{T},Z^{T}\rangle_{t}$ and $Y^{T}_{t}$ to the same limit. Our goal is to show that $\langle Z^{T},Z^{T}\rangle_{t}$ and $Y^{T}_{t}$ converge in probability to the same limit.
First, we establish a crucial identity. From the definitions of $C_{T}$ and $D_{T}$, we have the scaling factor for the martingale increments as $(C_{T}D_{T})$. Its square is:
$(C_{T}D_{T})^{2}=\left(\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}\right)\cdot\left(\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\right)^{2}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}=D_{T}.$
Using this, the predictable quadratic variation is:
$\langle Z^{T},Z^{T}\rangle_{t}=(C_{T}D_{T})^{2}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}=D_{T}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}.$
By the definition of $\Lambda_{t}^{T}$, this is precisely:
$\langle Z^{T},Z^{T}\rangle_{t}=\Lambda_{t}^{T}.$
Therefore, showing that $\langle Z^{T},Z^{T}\rangle_{t}$ and $Y^{T}_{t}$ converge to the same limit is equivalent to showing that $\Lambda_{t}^{T}$ and $Y^{T}_{t}$ converge to the same limit.
From Lemma 6, we have already proven that
$\sup_{t\in[0,1]}|Y_{t}^{T}-\Lambda_{t}^{T}|\to 0\quad\text{in probability}.$
This directly implies that the two processes converge to the same limit. Since the sequence $(Y^{T})$ is tight and converges to a limit process $Y$, it follows that $\Lambda_{t}^{T}\to Y_{t}$ in probability as well. Given our identity $\langle Z^{T},Z^{T}\rangle_{t}=\Lambda_{t}^{T}$, we can conclude that
$\langle Z^{T},Z^{T}\rangle_{t}\to Y_{t}\quad\text{in probability}.$
Combining (a) and (b), we have $Z^{T},Z^{T}]_{t}\to Y_{t}$ in probability.
3. Characterization of the Limit. Since $(Z^{T},Y^{T})$ converges in distribution to $(Z,Y)$, and $Z^{T},Z^{T}]_{t}\to Y_{t}$ in probability, it follows from the properties of weak convergence that the limit process satisfies $Y=[Z,Z]$.
Furthermore, we have established that $(Z^{T})$ is $\mathbb{C}$-tight. A sequence of martingales that is $\mathbb{C}$-tight and whose quadratic variation converges to a continuous process must converge to a continuous martingale (see, e.g., Proposition VI-3.26 in Jacod and Shiryaev (2013)). Since $Z^{T},Z^{T}]_{t}\to Y_{t}$ and $Y$ (as the limit of a $\mathbb{C}$-tight sequence $Y^{T}$) is continuous, the limit process $Z$ must be a continuous martingale.
5.2 Proof of Theorem 1
In fact, we only need to verify that the limit process satisfies equation (12). The proof of the Hölder property for $Y$ in Theorem 1 is exactly the same as Jaisson and Rosenbaum (2016). Like the continuous case in Jaisson and Rosenbaum (2016). We consider a converging subsequence $(Z^{T_{n}},Y^{T_{n}})$ and write $(Z,Y)$ its limit. Abusing notation slightly, we write $(Z^{T},Y^{T})$ instead of $(Z^{T_{n}},Y^{T_{n}})$. Using Skorohod’s representation theorem, there exists a probability space on which one can define copies in law of the $(Z^{T},Y^{T})$ converging almost surely under the skorohod topology to a random variable with the same law as $(Z,Y)$. We now work with this sequence of variables converging almost surely and their limit. The processes $Z$ and $Y$ being continuous, we obtain
$\sup_{t\in[0,1]}|Y_{t}^{T}-Y_{t}|\rightarrow 0,\ \sup_{t\in[0,1]}|Z_{t}^{T}-Z_{t}|\rightarrow 0.$
Let us now rewrite $\sum_{s=1}^{n}\lambda_{s}^{T}$. We can use Lemma 7 to obtain the expression of $\sum_{s=1}^{n}\lambda_{s}^{T}$, for all $n\in\mathbb{N}$, using Abel’s lemma, we obtain
$\sum_{s=1}^{n}\lambda_{s}^{T}=$ $n\mu^{T}+\sum_{s=1}^{n}\eta_{n-s}^{T}\sum_{u=1}^{s}\lambda_{u}^{T}+\sum_{s=1}^{n}\sum_{u=1}^{s-1}\eta_{s-u}^{T}(M_{u}^{T}-M_{u-1}^{T})$
$=$ $n\mu^{T}+\sum_{s=1}^{n}\eta_{n-s}^{T}\sum_{u=1}^{s}\lambda_{u}^{T}+\sum_{s=1}^{n}\eta_{n-s}^{T}M_{s}^{T}.$
Then we try to use Lemma 7, we first calculate
$\sum_{s=1}^{n-1}A_{n-s}^{T}\sum_{u=1}^{s}\eta_{s-u}^{T}M_{u}^{T}=$ $\sum_{u=1}^{n-1}M_{u}^{T}\sum_{s=1}^{n-u-1}A_{n-s-u}^{T}\eta_{s}^{T}=\sum_{u=1}^{n-1}M_{u}^{T}(A_{s-u}^{T}-\eta_{n-u}^{T}),$
where we use the fact that $A^{T}*\eta^{T}=A^{T}-\eta^{T}$. By Lemma 7, we obtain
$\sum_{s=1}^{n}\lambda_{s}^{T}=n\mu^{T}+\mu^{T}\sum_{s=1}^{n-1}sA_{n-s}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}M_{s}^{T}.$
Substituting $n$ with $\lfloor tT\rfloor$ and multiplying by $(1-a_{T})/(T^{\alpha}\nu^{*}\delta^{-1})$, denote
$u_{T}=\frac{\mu^{T}}{\nu^{*}\delta^{-1}T^{\alpha-1}},$
we can then decompose $\Lambda_{t}^{T}$ into three parts
$\Lambda_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=1}^{\lfloor tT\rfloor}\lambda_{s}^{T}=T_{1}+T_{2}+T_{3},$
where
$T_{1}=(1-a_{T})tu_{T},$
$T_{2}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\mu^{T}\sum_{s=1}^{\lfloor tT\rfloor-1}sA_{\lfloor tT\rfloor-s}^{T}=a_{T}u_{T}\sum_{s=1}^{\lfloor tT\rfloor-1}\frac{1}{T}\left(\frac{s}{T}\right)\rho^{T}\left(t-\frac{s}{T}\right),$
$T_{3}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor tT\rfloor-s}^{T}M_{s}^{T},$
and $\rho^{T}(\cdot)$ is defined in (10). Since $u_{T}$ converges to $1$, we get that $T_{1}\rightarrow 0$. By integrating by parts,
$T_{2}\sim$ $a_{T}u_{T}\int_{0}^{t}\rho^{T}(t-s)sds=a_{T}u_{T}\int_{0}^{t}F^{T}(t-s)ds$
$\rightarrow$ $\int_{0}^{t}F^{\alpha,\lambda}(t-s)ds=\int_{0}^{t}f^{\alpha,\lambda}(t-s)sds.$
We now turn to $T_{3}$. Remark that $M_{s}=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}Z_{s/T}^{T}$, so for sufficiently large $T$,
$T_{3}=$ $\frac{a_{T}}{\sqrt{T^{\alpha}(1-a_{T})\nu^{*}\delta^{-1}}}\sum_{s=1}^{\lfloor tT\rfloor-1}\frac{1}{T}\rho^{T}\left(t-\frac{s}{T}\right)Z_{s/T}^{T}$
$\sim$ $\frac{a_{T}}{\sqrt{T^{\alpha}(1-a_{T})\nu^{*}\delta^{-1}}}\int_{0}^{t}\rho^{T}(t-s)Z_{s}^{T}ds.$
Then we can imitate the Proof of Theorem 3.1 from Jaisson and Rosenbaum (2016), and obtain the same result, $T_{3}$ converges to
$\frac{1}{\nu^{*}\lambda}\int_{0}^{t}\lambda(t-s)^{\alpha-1}E_{\alpha,\alpha}(-\lambda(t-s)^{\alpha})Z_{s}ds.$
Since $Z$ is a continuous martingale, the fact that $Z_{t}=B_{Y_{t}}$ is a consequence of the Dambis-Dubin-Schwarz theorem: see, for example, Theorem V-1.65 in Revuz and Yor (2013).
5.3 Proof of Corollary 2
The proof extends the arguments used to establish Theorem 1. The core of the argument lies in analyzing the convergence of the scaled process $Y_{t}^{T}$ under the new time-dependent baseline intensity $\hat{\mu}_{T}(n)$.
The proof proceeds by decomposing the process $Y_{t}^{T}$ into its deterministic mean, $\mathbb{E}[Y_{t}^{T}]$, and its centered stochastic part, $Y_{t}^{T}-\mathbb{E}[Y_{t}^{T}]$, and then analyzing the limit of each component separately.
Let us denote the deterministic mean as $\bar{\Lambda}_{t}^{T}:=\mathbb{E}[Y_{t}^{T}]$. Using the expression for $\mathbb{E}[N_{n}^{T}]$ from Lemma 9 and applying the scaling factor, we can decompose $\bar{\Lambda}_{t}^{T}$ into several terms. After substituting the definition of $\hat{\mu}_{T}(n)$, the expression for $\mathbb{E}[N_{n}^{T}]$ becomes:
$\mathbb{E}[N_{n}^{T}]=\mu^{T}n+(1-\xi)\mu^{T}\sum_{s=1}^{n-1}sA_{n-s}^{T}+\xi\mu^{T}\left(\frac{n}{1-a_{T}}-\sum_{s=1}^{n-1}sA_{n-s}^{T}\right).$
Scaling this expression gives the components of the deterministic mean $\bar{\Lambda}_{t}^{T}$:
  • A term that vanishes in the limit:
    $T_{1}=\frac{(1-a_{T})\mu^{T}\lfloor Tt\rfloor}{T^{\alpha}\nu^{*}\delta^{-1}}.$
  • The main deterministic term that converges to a non-trivial limit:
    $T_{2}=\frac{(1-a_{T})\mu^{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\left[(1-\xi)\sum_{s=1}^{\lfloor Tt\rfloor-1}sA_{\lfloor Tt\rfloor-s}^{T}+\xi\left(\frac{\lfloor Tt\rfloor}{1-a_{T}}-\sum_{s=1}^{\lfloor Tt\rfloor-1}sA_{\lfloor Tt\rfloor-s}^{T}\right)\right].$
The centered stochastic part, $Y_{t}^{T}-\bar{\Lambda}_{t}^{T}$, is obtained by scaling the expression from Equation (20) in Lemma 9. We denote this scaled stochastic part as $T_{3}^{\text{stoch}}$:
  • The scaled centered stochastic part:
    $T_{3}^{\text{stoch}}:=Y_{t}^{T}-\bar{\Lambda}_{t}^{T}=\frac{1-a_{T}}{T^{\alpha}\nu^{*}\delta^{-1}}\left(M_{\lfloor Tt\rfloor}^{T}+\sum_{s=1}^{\lfloor Tt\rfloor-1}A_{\lfloor Tt\rfloor-s}^{T}M_{s}^{T}\right).$
Next, we analyze the limits of these terms as $T\to\infty$:
  • By Assumption 2, it is straightforward to show that $\lim_{T\to\infty}T_{1}=0$.
  • For $T_{2}$, we use the same Riemann sum approximation argument as in the proof of Theorem 1, where the sum $\sum_{s=1}^{n-1}sA_{n-s}^{T}$ corresponds to the integral of the CDF. In the limit, the sum of the deterministic parts converges:
    $\lim_{T\to\infty}\bar{\Lambda}_{t}^{T}=\lim_{T\to\infty}(T_{1}+T_{2})=(1-\xi)\int_{0}^{t}F^{\alpha,\lambda}(u)du+\xi\left(t-\int_{0}^{t}F^{\alpha,\lambda}(u)du\right).$
  • The convergence of the scaled stochastic part, $T_{3}^{\text{stoch}}$, was established in the proof of Theorem 1, as the structure of the martingale part is identical. Its limit is given by:
    $\lim_{T\to\infty}T_{3}^{\text{stoch}}=\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)B_{Y_{s}}ds.$
Combining these limits, the limit process $Y_{t}=\lim_{T\to\infty}Y_{t}^{T}$ is given by the sum of the limits of its deterministic and stochastic parts:
$Y_{t}=(1-\xi)\int_{0}^{t}F^{\alpha,\lambda}(u)du+\xi\left(t-\int_{0}^{t}F^{\alpha,\lambda}(u)du\right)+\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)B_{Y_{s}}ds.$
The process $\dot{Y}_{t}$ is the derivative of $Y_{t}$. Differentiating the deterministic part with respect to $t$ gives:
$\frac{d}{dt}\left(\lim_{T\to\infty}\bar{\Lambda}_{t}^{T}\right)=(1-\xi)F^{\alpha,\lambda}(t)+\xi\left(1-F^{\alpha,\lambda}(t)\right)=F^{\alpha,\lambda}(t)+\xi\left(1-F^{\alpha,\lambda}(t)\right).$
The derivative of the stochastic part, as established in Corollary 1, yields the stochastic integral term. Summing these derivatives gives the full stochastic Volterra equation for $\dot{Y}_{t}$ as stated in (21).
Finally, to confirm the initial condition, we evaluate at $t=0$. Since $F^{\alpha,\lambda}(0)=0$ and the stochastic integral is zero at $t=0$, we have:
$\dot{Y}_{0}=F^{\alpha,\lambda}(0)+\xi\left(1-F^{\alpha,\lambda}(0)\right)+0=\xi.$
This concludes the proof.
6 Future Work
In Section 2.1, we roughly used the scaling limit of the cumulative INAR($\infty$) process as an approximation of the integrated fractional CIR process. However, theoretically, the convergence rate of this scaling limit and the error analysis when using it as an approximation for the integrated fractional CIR process are still unknown. We will address this issue in future work. Another crucial direction for future research is the statistical estimation of the model’s parameters. While the present paper focuses on the theoretical scaling limit, the practical application of our INAR($\infty$) framework requires robust methods for estimating the key parameters, such as those governing the heavy-tailed kernel, from observed data. This estimation problem is non-trivial due to the long-memory nature of the process. A comprehensive treatment of this topic, including the development and analysis of a Least-Square Estimation (LSE) procedure for this class of models, is provided by the authors in the separate work Wang et al. (2024).
Acknowledgement
The authors would like to thank the Editor, the Associate Editor, and two anonymous referees for their insightful and constructive feedback, which led to a substantial improvement in the quality and clarity of this paper. We are particularly indebted to the first referee for an exceptionally thorough and meticulous review; their detailed comments on our mathematical arguments were invaluable in helping us to significantly strengthen the rigor of our proofs. We are also deeply grateful to the second referee for their insightful, high-level comments, which prompted us to clarify the novelty and broader contribution of our work. The authors are also grateful to Wei Xu, Yinhao Wu and Lingjiong Zhu for helpful comments.
Declarations
  • Conflict of interest/Competing interests: The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
  • Code availability: The code used in this study is provided as supplementary material. It is available under the MIT License and can be freely accessed and used for non-commercial purposes.
  • Author contribution: Yingli Wang led the research, authored the main manuscript, provided the majority of the proofs, prepared Figure 1, and developed Algorithm 1. Qinghua Wang contributed to the proof of Lemma 4 and played a key role in restructuring the manuscript. Ping He and Chunhao Cai supervised the project, provided critical revisions, and ensured the overall quality and coherence of the manuscript.
Funding Declarations
Yingli Wang is supported by the Fundamental Research Funds for the Central Universities in Shanghai University of Finance and Economics CXJJ2023-397.